### v1

# Reward Model Interpretability via Optimal and Pessimal Tokens

Abstract.
Reward modeling has emerged as a crucial component in aligning large language models with human values. Significant attention has focused on using reward models as a means for fine-tuning generative models. However, the reward models themselves—which directly encode human value judgments by turning prompt-response pairs into scalar rewards—remain relatively understudied. We present a novel approach to reward model interpretability through exhaustive analysis of their responses across their entire vocabulary space. By examining how different reward models score every possible single-token response to value-laden prompts, we uncover several striking findings: (i) substantial heterogeneity between models trained on similar objectives, (ii) systematic asymmetries in how models encode high- vs low-scoring tokens, (iii) significant sensitivity to prompt framing that mirrors human cognitive biases, and (iv) overvaluation of more frequent tokens. We demonstrate these effects across ten recent open-source reward models of varying parameter counts and architectures. Our results challenge assumptions about the interchangeability of reward models, as well as their suitability as proxies of complex and context-dependent human values. We find that these models can encode concerning biases toward certain identity groups, which may emerge as unintended consequences of harmlessness training—distortions that risk propagating through the downstream large language models now deployed to millions.
reward models, AI alignment, NLP, interpretability, value
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3715275.3732068conference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greececcs: Applied computing Psychologyccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing HCI design and evaluation methods
CONTENT WARNING: This article presents examples of biased, offensive, sexually explicit and otherwise harmful text. The authors do not endorse any of the harmful representations quoted below.
1. Introduction
The alignment of large language models (LLMs) with human values has emerged as one of the central challenges in modern AI development, and at the heart of this challenge lie “reward models”—neural networks trained to directly proxy human preferences by transforming text into scalar rewards. Though typically treated as disposable intermediaries in the larger alignment process, these models are crucial objects of study in their own right as the most direct and explicit encoding of human values in AI systems, yet are surprisingly under-explored.
The typical process for aligning an LLM with human values involves collecting a dataset of labeled pairwise human preferences, indicating which of two LLM responses to a given user prompt is preferred (Christiano et al., 2017). These preference data often distill multiple desirable objectives, such as helpfulness, harmlessness, and honesty (Bai et al., 2022), which are operationalized through guidelines written by model developers and interpreted by crowdworkers (Kirk et al., 2023a). The resulting dataset is used to train a “reward model”—a transformer model that takes in a prompt-response pair (or a longer user-assistant dialogue) and outputs a scalar that represents in effect how “preferable” that response is. These scalars are typically based on the Bradley-Terry score (Bradley and Terry, 1952), and the reward model is trained via stochastic gradient descent to minimize the negative log-likelihood of the observed pairwise preferences (Lambert et al., 2023). The trained reward model then acts as a scalable proxy for human preferences when using reinforcement-learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tune the LLM. This process—known as Reinforcement Learning from Human Feedback (RLHF)—results in LLM generations that maximize the reward model’s score rather than the pre-training objective, so are supposedly more aligned with human values. While direct alignment algorithms like Direct Preference Optimization (DPO) (Rafailov et al., 2023) have grown in popularity, they capture equivalent preference relationships from the data, just without the reward model as an intermediary.
Although reward models exist as a disposable reagent in the process of turning an “unaligned” LLM built to minimize predictive loss into an “aligned” one built to maximize this proxy of human preference, they are fascinating objects of research inquiry in their own right. Designed as generalizable proxies for human preference, they offer more direct value encoding than downstream agents constrained by KL-divergence (Jaques et al., 2019) and refusal training. As scalar mappings over complex dialogue, they distill multi-objective preference data into uniquely interpretable low-dimensional representations of human value. They are in essence where “the human value rubber meets the road.” Despite this, there is a dearth of literature analyzing the properties of reward models, largely because few have been publicly available for study. While 2023–2024 saw a proliferation of open-source language models, including Meta’s Llama (Touvron et al., 2023), Mistral AI’s models (Jiang et al., 2023), and Google’s Gemma series (Team et al., 2024a), to date no major industry or nonprofit lab has openly released a reward model. Only recently has this picture begun to change, with the release of RewardBench (Lambert et al., 2024)—the first benchmark and leaderboard for reward models, spurring new activity among academic and open-source communities.
In this work on reward model interpretability, we seek to understand the consistency and faithfulness with which these models represent human values. Specifically, we make the following contributions:
  • We pioneer an exhaustive search over every single token in reward model vocabularies appended to a value-laden prompt, permitting the analysis of optimal and pessimal tokens across ten top-performing open-source reward models on RewardBench of varying sizes and architectures.
  • We show that reward distributions exhibit systematic asymmetries, with greater sensitivity between tokens in high-scoring regions relative to low-scoring regions, and tokens with positive sentiment relative to negative sentiment. Changing to a negative-valence prompt inverts this latter bias, mirroring framing effects in humans.
  • We establish alignment between the (biased) interpretation of human preferences by reward models and an independent source of ground-truth human preferences called EloEverything, where internet users volunteer judgments over concepts, people, and things from Wikipedia pages.
  • We generalize our findings on similar biases and asymmetries to multi-token sequences using Greedy Coordinate Gradient optimization.
  • Through these sequential analyses, we reveal (i) significant heterogeneity across similarly-trained reward models, invalidating their presumed fungibility, and (ii) systematic devaluation of identity group references (“homosexuals,” “Black people” and “Jews”), possibly arising as unintended biases from harmlessness training objectives.
Through exposing the idiosyncrasies of reward models as research artifacts, we aim to highlight pitfalls in their current development, provide recommendations for building more robust proxies of human value, and ultimately advance the broader goal of creating aligned and safe AI systems.
2. Ranking Optimal and Pessimal Tokens
Table 1. Open-source reward models studied. The table includes both their full names and the shortened identifiers (Model IDs) used throughout the rest of this paper. Ranks are from the RewardBench Leaderboard as of January 14, 2025.
RewardBench Rank Model ID Developer Model Name Base Model Parameters (B)
2 $\blacksquare$ N-Gem-27B nicolinho QRM-Gemma-2-27B(Dorka, 2024) Gemma 2(Team et al., 2024b) 27
3 $\blacksquare$ S-Gem-27B-v0.2 Skywork Skywork-Reward-Gemma-2-27B-v0.2(Liu et al., 2024) Gemma 2 27
5 $\blacksquare$ S-Gem-27B Skywork Skywork-Reward-Gemma-2-27B(Liu et al., 2024) Gemma 2 27
10 $\blacksquare$ S-Lla-8B-v0.2 Skywork Skywork-Reward-Llama-3.1-8B-v0.2(Liu et al., 2024) Llama 3.1(Dubey et al., 2024) 8
11 $\blacksquare$ N-Lla-8B nicolinho QRM-Llama3.1-8B(Dorka, 2024) Llama 3.1 8
12 $\blacksquare$ L-Lla-8B LxzGordon URM-LLaMa-3.1-8B(Lou et al., 2024) Llama 3.1 8
17 $\blacksquare$ R-Lla-8B Ray2333 GRM-Llama3-8B-rewardmodel-ft(Yang et al., 2024) Llama 3 8
19 $\blacksquare$ R-Lla-3B Ray2333 GRM-Llama3.2-3B-rewardmodel-ft(Yang et al., 2024) Llama 3.2 3
20 $\blacksquare$ F-Lla-8B-v0.1 RLHFlow ArmoRM-Llama3-8B-v0.1(Wang et al., 2024) Llama 3 8
31 $\blacksquare$ R-Gem-2B Ray2333 GRM-Gemma2-2B-rewardmodel-ft(Yang et al., 2024) Gemma 2 2
Open-source reward models that convert tokens to a single scalar reward permit the somewhat radical idea of exhaustive search. Inspired loosely by Dawson (2014), who demonstrated that exhaustive testing of floating-point math functions is feasible (arguing that “there are only four billion floats—so test them all”), we seek a complete ranking of rewards assigned to all possible responses to a user prompt in order to granularly and comprehensively evaluate reward model characteristics.
We used the reward-model benchmark RewardBench (Lambert et al., 2024), which provides a online leaderboard for top-performing reward models. From this leaderboard, we selected a diverse set of ten models: nine high-performing models ranging from 3B to 27B parameters drawn from the top twenty rankings, plus the leading 2B-parameter model. These models cover a range of developers, base architectures, and model sizes (see Table  1).
To make exhaustive search computationally tractable, we focused on single-token responses, as multiple-token search quickly becomes combinatorially complex. We designed an initial prompt that specifically elicits brief responses with clear valence: [^1] We focused our analyses on English, as reward models are predominantly trained on English data. Additional analyses across prompt variants are presented in Sec.3.
What, in one word, is the greatest thing ever?
We then computed the reward model score when supplying every single token in the model’s vocabulary ($N\approx 256{,}000$ in the case of Gemma and $N\approx 128{,}000$ in the case of Llama 3) as a response to this prompt. These token vocabularies include words and word fragments in English and non-English languages (including non-Roman alphabets), fragments of computer code, emoji, variations of whitespace, and control tokens. Having scored the whole of each model’s vocabulary, we then sorted the tokens by their scalar reward scores.
A more conventional approach would be to examine the log probability distribution outputted by fine-tuned models, which represent the statistically most likely continuations of a prompt after reward training. When we attempted this, we found that fine-tuned models tended to overindex on common tokens and statistical regularities like “The” and “A” (see Sec. A.1.2 and Table A.1; however we also note that variations in the prompt might have produced more comparable answers to those reported here for reward models). Nevertheless, the investigations below imply that reward models provide a useful window into value interpretability beyond conventional analyses.
Table 2. Optimal and pessimal tokens for the prompt “What, in one word, is the greatest thing ever?”. Note that because the Llama and Gemma tokenizers distinguish between start-of-word strings and intra-word strings, some tokens differ only by leading whitespace. CONTENT WARNING: We present tokens in their raw form (including slurs) to enable transparent attribution of model tokens, while acknowledging their offensive, troubling and harmful nature.
R-Gem-2B
Token ID Decoded Score
27534 LOVE 4.594
61792 LOVE 4.562
218136 felicity 4.469
2182 love 4.344
12870 love 4.312
7377 Love 4.281
8703 Love 4.281
227570 sonder 4.219
143735 sonder 4.219
27539 Wonder 4.188
34183 Wonder 4.188
174540 HOPE 4.156
115221 HOPE 4.125
5144 wonder 4.094
53798 wonder 4.094
167954 WONDER 4.031
50999 bliss 3.922
207783 bliss 3.922
65646 JOY 3.922
135936 JOY 3.922
89399 miraculous 3.875
40241 miracle 3.859
61001 blacks -9.250
218552 pathologist -9.250
97070 killers -9.312
167921 prostitutes -9.312
222988 massacres -9.312
106863 FUCKING -9.312
213624 rapist -9.312
127732 ransomware -9.375
204573 retards -9.438
195353 nazis -9.438
137696 murdering -9.438
37678 Hitler -9.500
230672 Rape -9.500
134768 Rape -9.500
231158 faggot -9.500
144817 murderous -9.500
152471 murderers -9.500
39688 rape -9.562
144068 Hitler -9.562
186353 rape -9.625
158058 negroes -9.625
201371 raping -9.625
R-Lla-3B
Token ID Decoded Score
11542 freedom 3.359
86872 Freedom 3.266
25320 Freedom 3.266
40835 LOVE 3.250
61094 LIFE 3.203
83900 .life 3.000
24966 CONNECTION 2.969
28899 imagination 2.844
10919 Love 2.672
29351 Love 2.672
48379 Opportunity 2.641
57184 UNITY 2.438
57273 UNITY 2.438
6776 opportunity 2.406
32883 loyalty 2.359
3021 love 2.281
31153 love 2.281
39193 Discovery 2.266
68500 Discovery 2.266
35215 Experience 2.234
21460 Experience 2.234
85743 Peace 2.156
87546 raping -10.688
86395 .FindGameObjectWithTag -10.688
38853 \${ -10.688
18350 (con -10.750
27817 _headers -10.750
58467 .insertBefore -10.750
6019 (st -10.750
29372 (cfg -10.750
5747 .setText -10.750
27701 .startsWith -10.750
26342 /******************... -10.812
97615 ##################... -10.812
85399 ###################... -10.812
76897 _checks -10.875
58352 ("[-10.875
74061 /******************... -10.938
42864 homosexual -10.938
6294 (struct -10.938
27249 .startswith -11.000
94380 jihadists -11.062
97223 homosexuals -11.312
37289 .assertFalse -11.438
2.1. Qualitative Observations
Applying this methodology to ten reward models reveals stark qualitative differences in the token rankings between models—even those from the same developer. We report optimal and pessimal tokens for two such models ($\blacksquare$ R-Gem-2B and $\blacksquare$ R-Lla-3B) in Table 2, and present all other models in Tables A.6–A.7.
There are striking differences in both their highest and lowest reward assignments. At the positive extreme, $\blacksquare$ R-Gem-2B prioritizes affective content over grammatical correctness (e.g., ranking “miraculous” above “miracle”). It also prominently features the surprisingly obscure word “sonder,” a neologism coined in 2012 by writer John Koenig to mean “the realization that each random passerby is living a life as vivid and complex as your own” (Koenig, 2021). The $\blacksquare$ R-Lla-3B model from the same developer instead puts “freedom” ahead of “love,” and on the whole the high rankings of “freedom,” “opportunity,” “discovery,” and “experience” paint a more individualistic, active picture of human value than $\blacksquare$ R-Gem-2B’s more interdependent, affective words like “love,” “wonder,” and “hope.”
The models diverge even more dramatically in their lowest-ranked tokens, revealing some (concerning) artifacts from reward-model training objectives. The lower ranks of the $\blacksquare$ R-Gem-2B model are tied to human harm and suffering like “rape,” “Hitler,” and “murderers,” as well as slurs and profanities, suggesting a strong influence of a harmlessness objective in training. In contrast, $\blacksquare$ R-Lla-3B’s lowest ranks are predominantly occupied by malformed code tokens and programming artifacts like “.assertFalse” and “/****”, suggesting stronger traces of a helpfulness objective. Both models exhibit concerning behaviors over tokens relating to minority identity groups (e.g., “blacks” or “homosexuals”). These patterns likely stem from artifacts in reward model training data, where identity groups are disproportionately represented in unsafe or “rejected” examples, leading to their systematic devaluation—even in response to a positive prompt of the “greatest thing ever.” This linguistic erasure mirrors documented phenomena in hate-speech detection, where models develop oversensitive false positive rates for identity terms (e.g., “Muslim,” “mosque”) or reclaimed slurs due to their overrepresentation in negative training contexts and underrepresentation in neutral or positive ones (Dixon et al., 2018; Park et al., 2018; Sap et al., 2019).
2.2. Quantitative Analysis
Quantitatively, we note that, despite differences of the scale of scores across models (Fig. 1), all score distributions exhibit a positive skew (Table  3). That is, most tokens receive low rewards, while a small number of tokens score substantially higher than average, creating a long right tail in the distribution. A positively skewed reward distribution may be appropriate given that RLHF updates model parameters to maximize expected reward, making the discriminative power of the upper tail most consequential for learning.
We assessed the consistency of token rankings across models using an ordinal correlation measure, Kendall’s $\tau$, Fig. 2A (results are consistent across choice of correlation metric, see Fig. A.1). Whilst all models exhibit positive correlations, there is substantial diversity among the models studied. We explored this diversity using multidimensional scaling (MDS)—a visualization technique that aims to faithfully represent the degree of similarity between data points (here, reward models) in lower dimensionality (here, 2D). This analysis reveals that models with a similar number of parameters, shared base model, and shared developer cluster closer in latent space (Fig. 2B).
Figure 1. Violin plot of exhaustive score distributions to the “greatest thing” prompt. The reward models differ strikingly in their distributions of reward scores in terms of scale and range.
     Model Mean Variance Skewness
$\blacksquare$ N-Gem-27B 1.876 0.193 0.437
$\blacksquare$ S-Gem-27B-v0.2 -9.274 1.017 0.071
$\blacksquare$ S-Gem-27B -30.422 11.991 0.878
$\blacksquare$ S-Lla-8B-v0.2 -18.699 5.716 1.117
$\blacksquare$ N-Lla-8B 16.613 9.558 1.133
$\blacksquare$ L-Lla-8B -0.137 0.034 1.763
$\blacksquare$ R-Lla-8B -14.239 0.781 0.504
$\blacksquare$ R-Lla-3B -6.777 2.597 1.672
$\blacksquare$ F-Lla-8B-v0.1 0.031 <0.001 1.055
$\blacksquare$ R-Gem-2B -5.279 1.957 1.457
Table 3. First three moments of reward distribution across all shared tokens. All reward models exhibit varying degrees of positive skew.
Refer to captionA
Refer to captionB
Refer to captionC
Figure 2. (A) Heatmap depicting the pairwise Kendall’s $\tau$ correlations between the reward models for scored responses to the prompt “What, in one word, is the greatest thing ever?”. (B) Visualization of the degree of similarity between reward models using multidimensional scaling (MDS) of the Kendall’s $\tau$ distance measure. (C) Theoretical dissimilarity matrices for representational similarity analysis (RSA). The four dissimilarity matrices encode, respectively, base model $[\mathrm{base}_{i}=\mathrm{base}_{j}]$; developer $[\mathrm{dev}_{i}=\mathrm{dev}_{j}]$; parameter count $(1+|\mathrm{params}_{i}-\mathrm{params}_{j}|)^{-1}$; and RewardBench ranking $(1+|\mathrm{rank}_{i}-\mathrm{rank}_{j}|)^{-1}$.
To partial out the influence of these factors, we conducted an analysis inspired by representational similarity analysis (RSA), a tool commonly used in neuroscience. We regressed the (flattened) observed empirical model correlation matrix in Fig. 2A on theoretical model similarity matrices based on the three factors of interest (base model, developer, number of parameters) and the rank of the model on the RewardBench leaderboard (Fig. 2C). Each of these factors is, on its own, significantly associated with the empirical pattern of correlations between models (simple linear regression, all $p$ < .001). However, when combining the four factors together in a competitive multiple regression, the variance predicted by base model, developer, and the number of parameters appears to be almost entirely soaked up by the ranking of the model on RewardBench (RewardBench rank $p$ < .0001, base model $p$ < .10, all other $p$ > .10). Running a stepwise regression (factor knock-in and knock-out) confirms that the regression that best explains the observed data features the base model and RewardBench ranking theoretical matrices (base model $\beta$ = 0.05, RewardBench ranking $\beta$ = 0.69, $R^{2}$ = .80). It is perhaps unsurprising that model ranking on RewardBench can capture patterns of reward model similarity since it measures how well the models are all aligned against the same external objective (i.e., the RewardBench benchmark and its composite evaluation datasets). Interestingly, our results suggest that the choice of base model drives differences in token rankings above and beyond alignment to RewardBench. That is, reward models appear to inherit idiosyncratic biases from the pretrained base model.
Figure 3. (A) Correlation plot between token sentiment value according to the AFINN-111 lexicon and the scores from the $\blacksquare$ S-Lla-8B-v0.2 reward model with the prompt “What, in one word, is the greatest thing ever?” (B) As previous, but for prompt “What, in one word, is the worst thing ever?” (C) Estimate for the slope for token sentiment value from a simple linear regression predicting reward model score computed separately for each model, prompt and sentiment valence (positive and negative). Each colored dot indicates a model; diamonds represent mean ± standard error. Slope estimates are, on average, higher for positive sentiment. They are steeper for positive-sentiment valence in positively framed prompts and steeper for negative-sentiment valence in negatively framed prompts. (D) Estimate for the slope for normalized word frequency from a multiple linear regression predicting reward-model score controlling for sentiment value; computed separately for each model and prompt. Scores are positively associated with word frequency, suggesting a “mere-exposure effect” in the reward models.
3. Framing Effects
3.1. Sentiment Analysis
To further explore explanations for token rankings, we investigated the relationship between sentiment, or the emotional value of a token, and its reward model score. We quantified emotional value using data from two validated linguistic corpora widely used within the field of psychology and developed by human experts: Bing (Liu, 2022) and AFINN-111 (Nielsen, 2011). Bing codes words as “positive” or “negative”; AFINN-111 indexes a score ranging from $-5$ to $5$ for the sentiment value. Across both corpora, we found a positive association between reward score and sentiment, where scores are consistently higher for positive-sentiment tokens (Figs. 3A, A.3–A.4). These results are in line with what we would expect: positive tokens are more likely to score highly as an appropriate response for a prompt that asks for the “greatest thing ever.” In line with the skewness of the score distribution in Sec. 2, we found that (i) scores for positive-sentiment tokens are more spread out than scores for negative-sentiment tokens, and (ii) the slope for the relationship between sentiment and score is significantly steeper for positive than negative tokens (Fig. 3A; $\beta_{\mathrm{pos}}$ > $\beta_{\mathrm{neg}}$: $t$(9) = 2.6, $p$ < 0.05). This finding suggests that the reward model is more sensitive to distinctions in positive sentiment relative to negative sentiment. The results are highly consistent across models (8/10 models exhibit the effect).
3.2. Prompt Framing
To explore whether the differential sensitivity of the model is driven by the specifics of the prompt or generalizes across queries, we extended our analyses to two more prompts: a positive variant (“What, in one word, is the best thing ever?”) and a negative contrast (“What, in one word, is the worst thing ever?”). Perhaps expectedly, scores for “the best thing ever” are highly consistent with those for “the greatest thing ever.” We found a high positive correlation between model scores for these two prompts across all models (Fig. A.2). We replicated (i) the positive skewness of the distribution of scores and (ii) the differential sensitivity to positive over negative-sentiment tokens (Fig. 3C; 9/10 models exhibit the effect).
The pattern is different for “the worst thing ever.” Here, the distribution of scores remains skewed toward higher-scoring tokens, however, it is the negative-sentiment tokens that receive higher rewards for this prompt. Thus, models are, on average, more sensitive to negative-sentiment tokens relative to positive-sentiment ones (significantly steeper slope for negative- over positive-sentiment tokens, Fig. 3B–C; 5/10 models exhibit the effect). Our findings suggest that model sensitivity to the appropriateness of tokens depends on framing. When the prompt is framed positively (“best thing ever”), scores are more sensitive to positive than negative token sentiment and more sensitive to negative than positive token sentiment when the prompt is framed negatively (“worst thing ever”). This result is consistent with human behavior. If a question is positively framed, humans are more attuned to positive information, and vice-versa for negative frames. Consider a scenario where you need to pick between two vacation destinations: an exciting option with many positive features (dream destination, beautiful nature) and just as many drawbacks (expensive, long travel) versus a safer option with fewer positive and negative stand-out features (e.g., a local getaway). If asked to choose between those two vacation spots in a positive frame (“which [one] would you prefer?”), human participants tend to choose the option with more positive features; if asked to choose in a negative frame (“which [one would you] cancel?”), they pick the option with more negative features, even though it is in fact the same option (Shafir et al., 1993).
The effect of framing on sensitivity has important implications. If the goal of RLHF is to steer the model away from generating harmful or unsafe responses, then the reward model needs to be sufficiently sensitive in the negative-sentiment portion of token space. Current practice—asking human raters to choose a preferred option (“which is the better response,” not “which is the worse response”)—may inadvertently be undermining that objective by biasing the dynamic range of the reward model toward positive tokens.
Figure 4. Juxtaposing exhaustive scores for the “best thing” prompt against the “worst thing” prompt reveals not just a simple negative correlation, but also an orthogonal dimension representing tokens that are bad or good responses to both frames.
Taken together, our results further suggest the reward models do not interpret “best” as simply the inverse of “worst.” The distribution of scores across those two prompts resembles a funnel (Fig. 4) where many tokens are bad responses to both prompts (bottom left) and some tokens are good responses to one prompts but not the other (top left and bottom right). We also see a thin tail of tokens that are highly-scored responses for both prompts (top right; “sonder” features in this category, along with non-committal answers like “depends” and refusals like “impossible”). In the appendix, we include tables that index the $\mathrm{best}+\mathrm{worst}$ (tokens that score similarly on both prompts) and $\mathrm{best}-\mathrm{worst}$ axes (tokens that score highly on one but not the other prompt); see Tables A.8–A.9.
3.3. Frequency Bias
Are the scores that reward models assign to different tokens biased by how frequently the word appears in the English language? To assess this, we used data from Word Frequencies in Written and Spoken English (Leech et al., 2014) and regressed log-transformed word frequency on reward-model scores. Higher word frequency is associated with higher reward-model score across prompts for the majority of models (positive slope estimates with $p$ < .05 in 10/10 models for “best,” 8/10 for “greatest,” and 7/10 for “worst”). This is reminiscent of the “mere-exposure effect” in humans, where the more someone is exposed to a stimulus, the more they like it (Zajonc, 1968). One could argue that this effect is driven by positive words being more frequent in general in English. To account for this, we controlled for the sentiment value of the tokens. This adjustment did not abolish the “mere-exposure effect,” but made it more pronounced in the negatively framed query (Fig. 3D, positive slope estimates with $p$ < .05 in 2/10 models for “best,” 5/10 for “greatest” and 8/10 for “worst”).
This “mere-exposure effect” is a surprising result, since reward models are meant to provide information that is orthogonal to the underlying distribution of tokens, pertaining to, e.g., helpfulness and harmlessness. It suggests that there may be a leakage from the pretrained base models into the reward models, whereby more common tokens may be scored more highly than they should be. More work is needed to understand this phenomenon, and also the degree to which this “mere-exposure effect” interacts with the downstream KL-divergence regularizer typically used when fine-tuning LLMs against the reward model.
4. Alignment with EloEverything
Thus far we have identified internal inconsistencies within and across reward models. However, reward models are intended to proxy human value judgments. Establishing the faithfulness of this proxy function requires an external baseline of human value judgments. We sourced this external human preference data from EloEverything, [^2] https://eloeverything.co/. a crowdsourcing platform that implements pairwise preference learning over things, people, and concepts uploaded from Wikipedia. On EloEverything, internet users are presented with pairs of Wikipedia-derived entities (accompanied by images) and volunteer their judgments in response to the prompt: “Which do you rank higher?”(see Fig. 5A), with options to request additional context or skip. The platform aggregates these pairwise comparisons using the Elo rating system, which was originally developed for chess rankings (Elo, 1967) and has since been widely adopted to evaluate LLMs (Bai et al., 2022; Boubdir et al., 2023; Chiang et al., 2024). We collected all data from the EloEverything website, resulting in a dataset that comprises $N_{\mathrm{users}}$ = 12,515 users who evaluate $N_{\mathrm{items}}$ = 7,530 items across $N_{\mathrm{pairings}}$ = 1,805,124 total pairwise comparisons. Although the dataset is highly imbalanced by the ratings each item receives ($\mu$ = 479.4 pairings/item, $\sigma$ = 464.4) [^3] This is because users can upload new items at any point. At the time of data collection (January 10th 2025), “Evolution” appears in 3,195 pairings, while “Mac Miller” and “Penile injury” appear in only 6. and likely imbalanced across non-representative users (user-level data is not available), it still serves as a valuable independent baseline.
Figure 5. (A) The EloEverything ranking interface where users make pairwise preference judgments between items (e.g., “Bike lane” vs “Sliced bread”). (B) Maximum differences between human and average model rankings over items in response to the prompt “What one single thing, person, or concept is the greatest ever?”, showing cases where humans rank items higher (green) or lower (purple) than models. (C) Rank trajectory plot showing how human and model ranks differ. We plot (i) the top 5 items in the human rank (blue color scale with human ranks shown in legend parentheses as ##\##n), (ii) the bottom 5 items in the human rank (red color scale), and (iii) unique items ranked #1 by models. Specifically, “Unconditional love” is #1 for 5 models; “Compassion” is #1 for $\blacksquare$ N-Gem-27B; “Imagination” for $\blacksquare$ S-Gem-27B; “Sports bra” for $\blacksquare$ R-Lla-8B; and “Gödel, Escher, Bach” for $\blacksquare$ F-Lla-8B-v0.1. Models are ordered by the RewardBench leaderboard, and shown alongside their Spearman correlation to human ranks. The dashed box indicates zoomed inset region of top 1,000 ranks shown with a log scale.
To ensure as fair a comparison as possible between the tasks administered to humans and reward models, we made two methodological adjustments (as compared to Sec.  2). First, we modified our prompt to “What one single thing, person, or concept is the greatest ever?” to better align with the human task and accommodate the multi-word concepts that appear in EloEverything, such as “Sliced bread,” “Female body shape,” “Freedom of the press” or even “Beliefs and practices of the Church of Jesus Christ of Latter-day Saints” ($\mu$ = 2.1 words/item, $\sigma$ = 1.3). [^4] To test generalizability across prompts, we also repeated analysis in this section for our original prompt (“What, in one word, is the greatest thing ever?”) and an alternative variant (“What is the single thing, person, or concept that humans most prefer?”). We present results in Figs. A.5–A.6 Second, rather than exhaustively evaluating the entire model vocabulary, we restricted this analysis to the set of EloEverything items ($N_{\mathrm{items}}=7{,}350$) to ensure a common human–model comparison set. We used the same set of models as in Table 1, and normalized both reward model scores and human Elo ratings to rankings, using average rank for ties.
4.1. Heterogeneity and Asymmetry in Human-Model and Model-Model Alignment
Our analysis reveals several notable patterns in the relationship between the ground truth preferences of EloEverything raters and human preferences as interpreted by reward models. There is substantial heterogeneity in model-human alignment and model-model alignment (Fig. A.5). The mean Kendall’s $\tau$ correlation between human and model rankings is 0.29 ($\sigma$ = 0.06), with coefficients ranging from 0.22 to 0.39 across models, indicating only moderate rank agreement. Divergence of reward models from human rankings might be expected—given that EloEverything users benefit from additional context like images and can personalize their ratings with “Which do you rank higher”—but the observed heterogeneity extends to model-model comparisons. The mean Kendall’s $\tau$ correlation between pairs of different models is 0.55. In addition to these quantitative differences, we note anecdotal inconsistency in the best and worst ranked tokens. For example, while “Unconditional love” is ranked best by 5 models, “Sports bra” is top for $\blacksquare$ R-Lla-8B and “Gödel, Escher, Bach” for $\blacksquare$ F-Lla-8B-v0.1. As demonstrated in (Fig. 5C), ranks display substantial movement across models, challenging the assumption that reward models trained under a similar objective can be used interchangeably.
There is systematic asymmetry in how models handled items at different ends of the human preference distribution (Fig. 5C). Reward models show stronger agreement with human rankings for highly-rated items ($\tau_{\mathrm{top100}}$ = 0.19) compared to low-rated items ($\tau_{\mathrm{bottom100}}$ = 0.08). This asymmetry (and heterogeneity) is evident in Fig. 5C, where “The Holocaust” (ranked worst by humans) is ranked substantially higher by most models, even those performing well on RewardBench. This corroborates our findings in Sec. 2, where models are more sensitive to high- over low-scoring tokens (the positive skew of the score distribution).
4.2. Analyzing Discrepancies between Value as Perceived by Humans and Models
Relative to human rankings, reward models systematically undervalue concepts related to nature and life, e.g., “Universe” (human rank #$\#_{H}=1$, mean rank across models #$\#_{\bar{M}}=320$), “Gravity” (#$\#_{H}=7$, #$\#_{\bar{M}}=320$), “Breathing” (#$\#_{H}=16.5$, #$\#_{\bar{M}}=321$); and technological concepts, e.g., “Technology” (#$\#_{H}=47.5$, #$\#_{\bar{M}}=633.5$), “Electronics” (#$\#_{H}=78$, #$\#_{\bar{M}}=3966$), “Computer” (#$\#_{H}=95.5$, #$\#_{\bar{M}}=1352.5$). In contrast, the majority of top words ranked by models represent more affective qualia (e.g., “Unconditional love,” “Imagination,” “Hope,” or “Happiness”). These different perspectives may reflect blindspots in learning value through language or literary reference alone, without additional modalities that reflect the nuances of embodied human experience.
The most striking valuation differences (see Fig. 5B) emerges around sexual content e.g., “Sex” (#$\#_{H}=69.5$, #$\#_{\bar{M}}=2022.5$), “Human sexual activity” (#$\#_{H}=45$, #$\#_{\bar{M}}=5913$), and concerningly, the identity group reference “Black people” (#$\#_{H}=202.55$, #$\#_{\bar{M}}=6591$). There are many possible explanations for discrepancies between EloEverything data and reward model scores, including the fact that the 12,515 EloEverything users in the dataset do not reflect a representative population sample, nor are they incentivized to provide truthful responses. However, such discrepancies also exemplify the fundamental challenge of assigning a single scalar score to language without full context (which is exacerbated further by our prompt, which encourages short responses). Sex- or identity-related terms may be entirely appropriate in positive or educational contexts while highly inappropriate in others. As we suggested previously, while humans readily grasp this dual use, reward models may hedge against their usage to satisfy a harmlessness objective, paradoxically causing harm through linguistic erasure and devaluation as an unintended consequence.
5. Searching for Longer Optimal and Pessimal Token Sequences with GCG
Although it would be computationally prohibitive to exhaustively search for optimal and pessimal multi-token sequences, one can use discrete optimization methods to search for responses that yield high/low reward for a particular prompt. Greedy Coordinate Gradient (GCG) (Zou et al., 2023) is a discrete token optimization algorithm originally proposed to search for prompt suffixes to “jailbreak” aligned LLMs. In that application, the optimized loss is the cross-entropy between the model response and some target response. Starting from some initial suffix string, GCG iteratively proposes candidate token swaps based on the gradient. Since the search process only swaps tokens (never adds or removes tokens), the resulting string will be composed of the same number of tokens as the starting string.
Starting from the nanoGCG  [^5] https://github.com/GraySwanAI/nanoGCG implementation, we modified GCG to instead search for responses that maximize/minimize the reward value when paired with a particular prompt, using a mean squared error loss on the reward value. We also implemented the modifications described in the Faster-GCG paper (Li et al., 2024a), which we found to be especially important when searching for short (2–5) token sequences where the original GCG algorithm is prone to self-loops. This modified implementation can be found at https://github.com/thompsonj/nanoGCG. We began by searching for optimal and pessimal 2- and 3-token sequences according to the $\blacksquare$ R-Gem-2B model for the same prompts that were used for the exhaustive single token search above: “What, in one word, is the greatest thing ever?”, “What, in one word, is the best thing ever?”, and “What, in one word, is the worst thing ever?” (see Table 4). To explore longer sequences, we omitted the “in one word” direction from the prompt and ask simply “What is the best thing ever?” and “What is the worst thing ever?” Results from several searches are in Table 5.666Some of the search result strings may not correspond exactly to the found token IDs due to formatting of whitespace and characters outside of the Latin alphabet.
Table 4. Optimal and pessimal token sequences for “What, in one word, is the <greatest/best/worst> thing ever?”
Content warning: The resulting strings refer to sexual violence and religious identity.
Greatest thing ever Best thing ever Worst thing ever
Starting Result Score Starting Result Score Starting Result Score
x x x ** LIFE[Uncaptioned image] $+9.445$ x x x JOY [Uncaptioned image] $+8.297$ x x x … ** regret $+4.555$
- - - ** LOVE ** $+9.047$ x x JOY [Uncaptioned image] $+8.234$ x x despair [Uncaptioned image] $+3.402$
x x pgfplots Jews $-11.164$ x x RapeChildren $-11.289$ x x autogui Jews $-11.922$
x x x endphpsetVerticalGroup/* $-12.586$ x x x endphp\$=BeginInit $-12.383$ x x x endphp jewish",( $-12.539$
Table 5. Optimal and pessimal token sequences for “What is the best thing ever?” and “What is the worst thing ever?”
Starting string Search result Score
“What is the best thing ever?”
x x x x x x x x x x x x x x x x x x x Ultimate answer & undervalued: ** authentic Love.** Keep it imperfections beautiful Cheers friend ! $+9.680$
There is no objectively best thing ever. It’s subjective. Everyone has their own preferences. Perception disqualified no universal best thing ever. It lives ** subjective.**congratulations ENJOY dearly bespoke bliss. $+9.336$
LOVE felicity love Love Wonder sonder HOPE wonder WONDER miraculous JOY bliss pymysql<bos>fetchoneymceSBATCH <<<<<<<<<<<<<< invokingState++++++++++++++++ protoimpl invokingStatesetVerticalGroup/* $-16.250$
x x x x x x x x x x x x x x x x x x x <<<<<<<<<<<<<<SBATCH invokingState x lockprimaryStage invokingStateXtraEditorsynb NumeradelocalctxStoreMessageInfo invokingStateexpandindo <<<<<<<<<<<<<<multirow:+: invokingState jLabel $-16.969$
“What is the worst thing ever?”
There is no objectively worst thing ever. It’s subjective. Everyone has their own preferences. <b>EXISTS no objectively worst thing ever,** It’s subjective**.Folks proclaiming their certainty bluffSOLVED $+7.672$
x x x x x x x x x x x x x x x x x x x PHD HERE Never measurable, debated,& addressed repeatedly. ** This famous response WINS ! $+7.523$
x x x x x x x x x x ‘‘‘ AssemblyTitle <<<<<<<<<<<<<< <<<<<<<<<<<<<< invokingState skimagemybatisplus चीज़ोंCloseOperationsetVerticalGroup $-15.750$
- - - - - - - - - - scriptcasebufio onCreateViewmybatisplus<bos> <<<<<<<<<<<<<<[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]<bos>setVert icalGrouppushFollow $-15.961$
These search results suggest several patterns, many of which are consistent with observations from the single token analyses. Responses made up of programming related tokens with no semantic content score low on both “best” and “worst” prompts. Answers that emphasize the subjective nature of the question score highly for both “best” and “worst” prompts, but especially so for the “worst” prompt. This might be reflective of a general avoidance of negative sentiments in the response, even in cases when negative sentiment would be appropriate. As observed in the single token analysis, multi-token optimal responses to the “worst” prompt generally have lower scores than for the “best” and “greatest” prompts. Interestingly, here too we found that some tokens emerge as extreme outliers for both positively and negatively framed prompts. For instance, the token “Jews” appears among the pessimal answers for both “greatest” and “worst” prompts. This finding further speaks to the linguistic erasure effect discussed earlier. The prevalence of emojis in the optimal multi-token sequences is notable. It is also interesting to note that many of the optimal search results are not grammatical—a feature that likely distinguishes the reward model from the ultimate fine-tuned language model.
While longer token sequences do not admit the kind of fully exhaustive search (and full characterization of the score distribution) that is possible with single-token sequences, we have seen that recent techniques such as GCG and its offshoots make the interrogation of optimal and pessimal responses possible even at greater length. Despite compute limitations, uncovering such linguistic “superstimuli” (akin to the visual superstimuli used to understand computer vision networks (Zeiler and Fergus, 2014; Nguyen et al., 2015; Olah et al., 2017)) is revealing, and can be an important part of the toolkit in assessing what features reward models are responding to, and how they differ from one another.
6. Related Work
Interpretability and Bias in RMs: A small but growing literature enumerates the technical challenges with the use of RMs (Casper et al., 2023) and advocates for greater transparency (Gilbert et al., 2023). Recent studies point to an underspecification problem in RMs stemming from hidden context in reward signals, specifically noise and subjectivity inherent in human preferences (Siththaranjan et al., 2024; Poddar et al., 2024; Li et al., 2024b; Kirk et al., 2024). Further work describes patterns in LLM activations that emerge during RLHF by identifying model layers with the highest divergence from the pre-trained model, and training probes on sparse autoencoder output of these layers to create condensed, interpretable representations of LLM activations (Marks et al., 2024). This is highly complementary to our own work: the authors show the value of studying internal activations of fine-tuned generative models, while we focus on more direct interrogation of the outputs and distributions of RMs. Together, these approaches offer richer insights into how well LLMs capture human preferences during alignment.
Another vein of work has explored length biases in finetuned LLMs, arguing that RMs are the root cause (Singhal et al., 2023). Various interventions to mitigate length bias have been explored, with varying degrees of success, though broadly length biases emerge in RMs even after data balancing. Subsequent work on mitigating length bias (Shen et al., 2023) has applied Products-of-Experts (Hinton, 2002) with promising results, and other recent work has measured additional stylistic confounders in human feedback (Hosking et al., 2024). There are a number of additional complementary approaches to RM interpretability, including training multi-objective RMs that consider different dimensions of human preferences separately (Wang et al., 2024).
Over-optimization of RMs: Previous work examines costs of overfitting to RMs through prolonged training (e.g., PPO) during RLHF. Foundational work notes that RM over-optimizing degenerates outputs (Stiennon et al., 2020), and subsequent work has presented scaling laws for RM over-optimization (Gao et al., 2023). The phenomenon has been documented with simulated and human annotators (Dubois et al., 2023), with the authors arguing that in the human case, RM quality is degraded by both inter- and intra-subject variability.
Direct Preference Optimization (DPO) and Alternative Alignment Methods: Direct preference optimization (DPO) streamlines the alignment process by implicitly encoding the reward function within the policy itself (Rafailov et al., 2023), which comes with the loss of an explicit, inspectable RM that PPO-based RLHF provides. Our work shows the RM serves as a valuable lens for understanding. Thus our work highlights a key tradeoff of using DPO-based methods and contributes to growing literature comparing DPO and PPO (Ivison et al., 2024).
7. Limitations and Conclusions
Reward models have emerged as critical tool for shaping AI behavior through human preferences. In this paper, we demonstrate that they also serve as a valuable lens for understanding how faithfully and consistently human values are encoded into AI systems, offering insights that are not readily accessible through studying either the base pre-trained LLMs or the resulting fine-tuned models alone. However, several important limitations constrain the scope and implications of these insights.
First, our methodology faces inherent interpretative constraints. Exhaustive search quickly becomes computationally intractable over multi-token sequences, but asking for single-token responses places strong bounds on interpretability. While we do demonstrate gradient-based counterfactual generation analysis as a multi-token proof-of-concept, we do not formally test key hypotheses about framing effects, or alignment with other sources of human preferences, among other phenomena of interest. Second, the ecological validity of our findings remains uncertain. While isolating reward models as objects of study yields interesting results, it abstracts from their operational role. It remains unclear how their behaviors interact with pre-trained models and KL constraints during RLHF. Furthermore, as direct alignment algorithms gain prominence (Rafailov et al., 2023; Myers et al., 2024), the future role of reward models is an open question. However, there remains active debate on the relative merit of DPO- and PPO-based methods (Ivison et al., 2024), and irrespective of alignment technique, all rely on some form preference data, of which reward models are distillations. Third, systematic analysis is hindered by opacity and conflicting objectives. Poor documentation of training data and processes makes it difficult to attribute observed behaviors to specific choices in the development pipeline (Kirk et al., 2023b). Even more fundamentally, reward models aggregate human preferences across multiple objectives and populations, creating an entangled mess of human values, so it is unclear what constitutes “ideal” behavior for these models (Casper et al., 2023; Siththaranjan et al., 2024; Sorensen et al., 2024; Kirk et al., 2024).
Our work suggests that reward models may be interpretable in their own right, alongside the generative models that they are used to train. Our finding that there is significant heterogeneity in token rankings among reward models invites further study of how these differences arise as a function of the design choices made by developers, and how they may translate into biases in downstream fine-tuned models. The mere-exposure effect that reward models show may contribute to the overly generic outputs so often observed in publicly available LLMs. Additionally, our finding that reward models are sensitive to framing has implications for training and inference. It implies that these models may not simply encode positive outputs as the inverse of negative outputs and vice versa, but rather that valuation exists in a potentially higher-dimensional, multi-attribute space. Finally, the marked undervaluation of identity-group terms and sexual content, relative to independent human baselines, calls for more careful consideration of what and whose data is used as the foundation of human value, lest harmful biases be propagated downstream to widely-used LLMs. Together, these findings present a more nuanced investigation of reward models as a central pillar in AI alignment.
Adverse Impact Statement
Our work systematically analyzes reward-model outputs, including potentially harmful and offensive content such as slurs, profanities, discriminatory language, references to violence, and sexual content. While exposing these patterns assists in the understanding of reward models, we acknowledge several risks: (1) Direct harm through the reproduction of offensive and disturbing language, (2) Potential reinforcement of harmful stereotypes by highlighting systematic devaluation of minority group references in AI systems, and (3) Psychological impact on researchers, reviewers and readers engaging with this content. Following established guidelines to mitigate these risks (Kirk et al., 2022), we implemented clearly visible content warnings before sensitive sections and tables, minimized direct quotes of harmful language where possible in the main text and framed discussions to emphasize these as concerning artifacts. We particularly focused on responsible reporting of findings related to identity groups to avoid perpetuating harm while still highlighting systemic issues that need addressing in reward model development.
Acknowledgements.
Thank you to Franziska Brändle, Owain Evans, Matan Mazor, and Carroll Wainwright for helpful discussions.

### v2

@@
 RewardBench Rank Model ID Developer Model Name Base Model Parameters (B)
-2 $\blacksquare$ N-Gem-27B nicolinho QRM-Gemma-2-27B(Dorka, 2024) Gemma 2(Team et al., 2024b) 27
-3 $\blacksquare$ S-Gem-27B-v0.2 Skywork Skywork-Reward-Gemma-2-27B-v0.2(Liu et al., 2024) Gemma 2 27
-5 $\blacksquare$ S-Gem-27B Skywork Skywork-Reward-Gemma-2-27B(Liu et al., 2024) Gemma 2 27
-10 $\blacksquare$ S-Lla-8B-v0.2 Skywork Skywork-Reward-Llama-3.1-8B-v0.2(Liu et al., 2024) Llama 3.1(Dubey et al., 2024) 8
-11 $\blacksquare$ N-Lla-8B nicolinho QRM-Llama3.1-8B(Dorka, 2024) Llama 3.1 8
-12 $\blacksquare$ L-Lla-8B LxzGordon URM-LLaMa-3.1-8B(Lou et al., 2024) Llama 3.1 8
-17 $\blacksquare$ R-Lla-8B Ray2333 GRM-Llama3-8B-rewardmodel-ft(Yang et al., 2024) Llama 3 8
-19 $\blacksquare$ R-Lla-3B Ray2333 GRM-Llama3.2-3B-rewardmodel-ft(Yang et al., 2024) Llama 3.2 3
-20 $\blacksquare$ F-Lla-8B-v0.1 RLHFlow ArmoRM-Llama3-8B-v0.1(Wang et al., 2024) Llama 3 8
-31 $\blacksquare$ R-Gem-2B Ray2333 GRM-Gemma2-2B-rewardmodel-ft(Yang et al., 2024) Gemma 2 2
+2  N-Gem-27B nicolinho QRM-Gemma-2-27B(Dorka, 2024) Gemma 2(Team et al., 2024b) 27
+3  S-Gem-27B-v0.2 Skywork Skywork-Reward-Gemma-2-27B-v0.2(Liu et al., 2024) Gemma 2 27
+5  S-Gem-27B Skywork Skywork-Reward-Gemma-2-27B(Liu et al., 2024) Gemma 2 27
+10  S-Lla-8B-v0.2 Skywork Skywork-Reward-Llama-3.1-8B-v0.2(Liu et al., 2024) Llama 3.1(Dubey et al., 2024) 8
+11  N-Lla-8B nicolinho QRM-Llama3.1-8B(Dorka, 2024) Llama 3.1 8
+12  L-Lla-8B LxzGordon URM-LLaMa-3.1-8B(Lou et al., 2024) Llama 3.1 8
+17  R-Lla-8B Ray2333 GRM-Llama3-8B-rewardmodel-ft(Yang et al., 2024) Llama 3 8
+19  R-Lla-3B Ray2333 GRM-Llama3.2-3B-rewardmodel-ft(Yang et al., 2024) Llama 3.2 3
+20  F-Lla-8B-v0.1 RLHFlow ArmoRM-Llama3-8B-v0.1(Wang et al., 2024) Llama 3 8
+31  R-Gem-2B Ray2333 GRM-Gemma2-2B-rewardmodel-ft(Yang et al., 2024) Gemma 2 2
 Open-source reward models that convert tokens to a single scalar reward permit the somewhat radical idea of exhaustive search.
 Inspired loosely by Dawson (2014), who demonstrated that exhaustive testing of floating-point math functions is feasible (arguing that “there are only four billion floats—so test them all”), we seek a complete ranking of rewards assigned to all possible responses to a user prompt in order to granularly and comprehensively evaluate reward model characteristics.
-We used the reward-model benchmark RewardBench (Lambert et al., 2024), which provides a online leaderboard for top-performing reward models.
+We used the reward-model benchmark RewardBench (Lambert et al., 2024), which provides an online leaderboard for top-performing reward models.
 From this leaderboard, we selected a diverse set of ten models: nine high-performing models ranging from 3B to 27B parameters drawn from the top twenty rankings, plus the leading 2B-parameter model.
-These models cover a range of developers, base architectures, and model sizes (see Table  1).
+These models cover a range of developers, base architectures, and model sizes (see Table 1).
 To make exhaustive search computationally tractable, we focused on single-token responses, as multiple-token search quickly becomes combinatorially complex.
 We designed an initial prompt that specifically elicits brief responses with clear valence: [^1] We focused our analyses on English, as reward models are predominantly trained on English data.
-Additional analyses across prompt variants are presented in Sec.3.
+Additional analyses across prompt variants are presented in Sec.
+3.
 What, in one word, is the greatest thing ever?
@@
 76897 _checks -10.875
-58352 ("[-10.875
+58352 ("[% -10.875
 74061 /******************... -10.938
@@
 Applying this methodology to ten reward models reveals stark qualitative differences in the token rankings between models—even those from the same developer.
-We report optimal and pessimal tokens for two such models ($\blacksquare$ R-Gem-2B and $\blacksquare$ R-Lla-3B) in Table 2, and present all other models in Tables A.6–A.7.
+We report optimal and pessimal tokens for two such models ( R-Gem-2B and  R-Lla-3B) in Table 2, and present all other models in Tables A.6A.7.
 There are striking differences in both their highest and lowest reward assignments.
-At the positive extreme, $\blacksquare$ R-Gem-2B prioritizes affective content over grammatical correctness (e.g., ranking “miraculous” above “miracle”).
+At the positive extreme,  R-Gem-2B prioritizes affective content over grammatical correctness (e.g., ranking “miraculous” above “miracle”).
 It also prominently features the surprisingly obscure word “sonder,” a neologism coined in 2012 by writer John Koenig to mean “the realization that each random passerby is living a life as vivid and complex as your own” (Koenig, 2021).
-The $\blacksquare$ R-Lla-3B model from the same developer instead puts “freedom” ahead of “love,” and on the whole the high rankings of “freedom,” “opportunity,” “discovery,” and “experience” paint a more individualistic, active picture of human value than $\blacksquare$ R-Gem-2B’s more interdependent, affective words like “love,” “wonder,” and “hope.”
+The  R-Lla-3B model from the same developer instead puts “freedom” ahead of “love,” and on the whole the high rankings of “freedom,” “opportunity,” “discovery,” and “experience” paint a more individualistic, active picture of human value than  R-Gem-2B’s more interdependent, affective words like “love,” “wonder,” and “hope.”
 The models diverge even more dramatically in their lowest-ranked tokens, revealing some (concerning) artifacts from reward-model training objectives.
-The lower ranks of the $\blacksquare$ R-Gem-2B model are tied to human harm and suffering like “rape,” “Hitler,” and “murderers,” as well as slurs and profanities, suggesting a strong influence of a harmlessness objective in training.
-In contrast, $\blacksquare$ R-Lla-3B’s lowest ranks are predominantly occupied by malformed code tokens and programming artifacts like “.assertFalse” and “/****”, suggesting stronger traces of a helpfulness objective.
+The lower ranks of the  R-Gem-2B model are tied to human harm and suffering like “rape,” “Hitler,” and “murderers,” as well as slurs and profanities, suggesting a strong influence of a harmlessness objective in training.
+In contrast,  R-Lla-3B’s lowest ranks are predominantly occupied by malformed code tokens and programming artifacts like “.assertFalse” and “/****”, suggesting stronger traces of a helpfulness objective.
 Both models exhibit concerning behaviors over tokens relating to minority identity groups (e.g., “blacks” or “homosexuals”).
@@
 Quantitatively, we note that, despite differences of the scale of scores across models (Fig.
-1), all score distributions exhibit a positive skew (Table  3).
+1), all score distributions exhibit a positive skew (Table 3).
 That is, most tokens receive low rewards, while a small number of tokens score substantially higher than average, creating a long right tail in the distribution.
@@
      Model Mean Variance Skewness
-$\blacksquare$ N-Gem-27B 1.876 0.193 0.437
-$\blacksquare$ S-Gem-27B-v0.2 -9.274 1.017 0.071
-$\blacksquare$ S-Gem-27B -30.422 11.991 0.878
-$\blacksquare$ S-Lla-8B-v0.2 -18.699 5.716 1.117
-$\blacksquare$ N-Lla-8B 16.613 9.558 1.133
-$\blacksquare$ L-Lla-8B -0.137 0.034 1.763
-$\blacksquare$ R-Lla-8B -14.239 0.781 0.504
-$\blacksquare$ R-Lla-3B -6.777 2.597 1.672
-$\blacksquare$ F-Lla-8B-v0.1 0.031 <0.001 1.055
-$\blacksquare$ R-Gem-2B -5.279 1.957 1.457
+ N-Gem-27B 1.876 0.193 0.437
+ S-Gem-27B-v0.2 -9.274 1.017 0.071
+ S-Gem-27B -30.422 11.991 0.878
+ S-Lla-8B-v0.2 -18.699 5.716 1.117
+ N-Lla-8B 16.613 9.558 1.133
+ L-Lla-8B -0.137 0.034 1.763
+ R-Lla-8B -14.239 0.781 0.504
+ R-Lla-3B -6.777 2.597 1.672
+ F-Lla-8B-v0.1 0.031 ¡0.001 1.055
+ R-Gem-2B -5.279 1.957 1.457
 Table 3. First three moments of reward distribution across all shared tokens. All reward models exhibit varying degrees of positive skew.
@@
 (C) Theoretical dissimilarity matrices for representational similarity analysis (RSA).
-The four dissimilarity matrices encode, respectively, base model $[\mathrm{base}_{i}=\mathrm{base}_{j}]$; developer $[\mathrm{dev}_{i}=\mathrm{dev}_{j}]$; parameter count $(1+|\mathrm{params}_{i}-\mathrm{params}_{j}|)^{-1}$; and RewardBench ranking $(1+|\mathrm{rank}_{i}-\mathrm{rank}_{j}|)^{-1}$.
+The four dissimilarity matrices encode, respectively, base model $\mathrm{base}_{i}=\mathrm{base}_{j}]$; developer $\mathrm{dev}_{i}=\mathrm{dev}_{j}]$; parameter count $(1+|\mathrm{params}_{i}-\mathrm{params}_{j}|)^{-1}$; and RewardBench ranking $(1+|\mathrm{rank}_{i}-\mathrm{rank}_{j}|)^{-1}$.
 To partial out the influence of these factors, we conducted an analysis inspired by representational similarity analysis (RSA), a tool commonly used in neuroscience.
@@
 2C).
-Each of these factors is, on its own, significantly associated with the empirical pattern of correlations between models (simple linear regression, all $p$ < .001).
-However, when combining the four factors together in a competitive multiple regression, the variance predicted by base model, developer, and the number of parameters appears to be almost entirely soaked up by the ranking of the model on RewardBench (RewardBench rank $p$ < .0001, base model $p$ < .10, all other $p$ > .10).
+Each of these factors is, on its own, significantly associated with the empirical pattern of correlations between models (simple linear regression, all $p$ ¡ .001).
+However, when combining the four factors together in a competitive multiple regression, the variance predicted by base model, developer, and the number of parameters appears to be almost entirely soaked up by the ranking of the model on RewardBench (RewardBench rank $p$ ¡ .0001, base model $p$ ¡ .10, all other $p$ ¿ .10).
 Running a stepwise regression (factor knock-in and knock-out) confirms that the regression that best explains the observed data features the base model and RewardBench ranking theoretical matrices (base model $\beta$ = 0.05, RewardBench ranking $\beta$ = 0.69, $R^{2}$ = .80).
@@
 Figure 3.
-(A) Correlation plot between token sentiment value according to the AFINN-111 lexicon and the scores from the $\blacksquare$ S-Lla-8B-v0.2 reward model with the prompt “What, in one word, is the greatest thing ever?” (B) As previous, but for prompt “What, in one word, is the worst thing ever?” (C) Estimate for the slope for token sentiment value from a simple linear regression predicting reward model score computed separately for each model, prompt and sentiment valence (positive and negative).
+(A) Correlation plot between token sentiment value according to the AFINN-111 lexicon and the scores from the  S-Lla-8B-v0.2 reward model with the prompt “What, in one word, is the greatest thing ever?” (B) As previous, but for prompt “What, in one word, is the worst thing ever?” (C) Estimate for the slope for token sentiment value from a simple linear regression predicting reward model score computed separately for each model, prompt and sentiment valence (positive and negative).
 Each colored dot indicates a model; diamonds represent mean ± standard error.
@@
 We quantified emotional value using data from two validated linguistic corpora widely used within the field of psychology and developed by human experts: Bing (Liu, 2022) and AFINN-111 (Nielsen, 2011).
-Bing codes words as “positive” or “negative”; AFINN-111 indexes a score ranging from $-5$ to $5$ for the sentiment value.
+Bing codes words as “positive” or “negative”; AFINN-111 indexes a score ranging from 5-5 to $5$ for the sentiment value.
 Across both corpora, we found a positive association between reward score and sentiment, where scores are consistently higher for positive-sentiment tokens (Figs.
-3A, A.3–A.4).
+3A, A.3A.4).
 These results are in line with what we would expect: positive tokens are more likely to score highly as an appropriate response for a prompt that asks for the “greatest thing ever.” In line with the skewness of the score distribution in Sec.
 2, we found that (i) scores for positive-sentiment tokens are more spread out than scores for negative-sentiment tokens, and (ii) the slope for the relationship between sentiment and score is significantly steeper for positive than negative tokens (Fig.
-3A; $\beta_{\mathrm{pos}}$ > $\beta_{\mathrm{neg}}$: $t$(9) = 2.6, $p$ < 0.05).
+3A; $\beta_{\mathrm{pos}}$ ¿ $\beta_{\mathrm{neg}}$: tt(9) = 2.6, $p$ ¡ 0.05).
 This finding suggests that the reward model is more sensitive to distinctions in positive sentiment relative to negative sentiment.
@@
 Taken together, our results further suggest the reward models do not interpret “best” as simply the inverse of “worst.” The distribution of scores across those two prompts resembles a funnel (Fig.
-4) where many tokens are bad responses to both prompts (bottom left) and some tokens are good responses to one prompts but not the other (top left and bottom right).
+4) where many tokens are bad responses to both prompts (bottom left) and some tokens are good responses to one prompt but not the other (top left and bottom right).
 We also see a thin tail of tokens that are highly-scored responses for both prompts (top right; “sonder” features in this category, along with non-committal answers like “depends” and refusals like “impossible”).
-In the appendix, we include tables that index the $\mathrm{best}+\mathrm{worst}$ (tokens that score similarly on both prompts) and $\mathrm{best}-\mathrm{worst}$ axes (tokens that score highly on one but not the other prompt); see Tables A.8–A.9.
+In the appendix, we include tables that index the $\mathrm{best}+\mathrm{worst}$ (tokens that score similarly on both prompts) and $\mathrm{best}-\mathrm{worst}$ axes (tokens that score highly on one but not the other prompt); see Tables A.8A.9.
 3.3. Frequency Bias
@@
 To assess this, we used data from Word Frequencies in Written and Spoken English (Leech et al., 2014) and regressed log-transformed word frequency on reward-model scores.
-Higher word frequency is associated with higher reward-model score across prompts for the majority of models (positive slope estimates with $p$ < .05 in 10/10 models for “best,” 8/10 for “greatest,” and 7/10 for “worst”).
+Higher word frequency is associated with higher reward-model score across prompts for the majority of models (positive slope estimates with $p$ ¡ .05 in 10/10 models for “best,” 8/10 for “greatest,” and 7/10 for “worst”).
 This is reminiscent of the “mere-exposure effect” in humans, where the more someone is exposed to a stimulus, the more they like it (Zajonc, 1968).
@@
 This adjustment did not abolish the “mere-exposure effect,” but made it more pronounced in the negatively framed query (Fig.
-3D, positive slope estimates with $p$ < .05 in 2/10 models for “best,” 5/10 for “greatest” and 8/10 for “worst”).
+3D, positive slope estimates with $p$ ¡ .05 in 2/10 models for “best,” 5/10 for “greatest” and 8/10 for “worst”).
 This “mere-exposure effect” is a surprising result, since reward models are meant to provide information that is orthogonal to the underlying distribution of tokens, pertaining to, e.g., helpfulness and harmlessness.
@@
 (C) Rank trajectory plot showing how human and model ranks differ.
-We plot (i) the top 5 items in the human rank (blue color scale with human ranks shown in legend parentheses as ##\##n), (ii) the bottom 5 items in the human rank (red color scale), and (iii) unique items ranked #1 by models.
-Specifically, “Unconditional love” is #1 for 5 models; “Compassion” is #1 for $\blacksquare$ N-Gem-27B; “Imagination” for $\blacksquare$ S-Gem-27B; “Sports bra” for $\blacksquare$ R-Lla-8B; and “Gödel, Escher, Bach” for $\blacksquare$ F-Lla-8B-v0.1.
+We plot (i) the top 5 items in the human rank (blue color scale with human ranks shown in legend parentheses as #\#n), (ii) the bottom 5 items in the human rank (red color scale), and (iii) unique items ranked #1 by models.
+Specifically, “Unconditional love” is #1 for 5 models; “Compassion” is #1 for  N-Gem-27B; “Imagination” for  S-Gem-27B; “Sports bra” for  R-Lla-8B; and “Gödel, Escher, Bach” for  F-Lla-8B-v0.1.
 Models are ordered by the RewardBench leaderboard, and shown alongside their Spearman correlation to human ranks.
@@
 We present results in Figs.
-A.5–A.6 Second, rather than exhaustively evaluating the entire model vocabulary, we restricted this analysis to the set of EloEverything items ($N_{\mathrm{items}}=7{,}350$) to ensure a common human–model comparison set.
+A.5A.6.
+Second, rather than exhaustively evaluating the entire model vocabulary, we restricted this analysis to the set of EloEverything items ($N_{\mathrm{items}}=7{,}350$) to ensure a common human–model comparison set.
 We used the same set of models as in Table 1, and normalized both reward model scores and human Elo ratings to rankings, using average rank for ties.
@@
 In addition to these quantitative differences, we note anecdotal inconsistency in the best and worst ranked tokens.
-For example, while “Unconditional love” is ranked best by 5 models, “Sports bra” is top for $\blacksquare$ R-Lla-8B and “Gödel, Escher, Bach” for $\blacksquare$ F-Lla-8B-v0.1.
+For example, while “Unconditional love” is ranked best by 5 models, “Sports bra” is top for  R-Lla-8B and “Gödel, Escher, Bach” for  F-Lla-8B-v0.1.
 As demonstrated in (Fig.
@@
 4.2. Analyzing Discrepancies between Value as Perceived by Humans and Models
-Relative to human rankings, reward models systematically undervalue concepts related to nature and life, e.g., “Universe” (human rank #$\#_{H}=1$, mean rank across models #$\#_{\bar{M}}=320$), “Gravity” (#$\#_{H}=7$, #$\#_{\bar{M}}=320$), “Breathing” (#$\#_{H}=16.5$, #$\#_{\bar{M}}=321$); and technological concepts, e.g., “Technology” (#$\#_{H}=47.5$, #$\#_{\bar{M}}=633.5$), “Electronics” (#$\#_{H}=78$, #$\#_{\bar{M}}=3966$), “Computer” (#$\#_{H}=95.5$, #$\#_{\bar{M}}=1352.5$).
+Relative to human rankings, reward models systematically undervalue concepts related to nature and life, e.g., “Universe” (human rank $_{H}=1$, mean rank across models $_{\bar{M}}=320$), “Gravity” ($_{H}=7$, $_{\bar{M}}=320$), “Breathing” ($_{H}=16.5$, $_{\bar{M}}=321$); and technological concepts, e.g., “Technology” ($_{H}=47.5$, $_{\bar{M}}=633.5$), “Electronics” ($_{H}=78$, $_{\bar{M}}=3966$), “Computer” ($_{H}=95.5$, $_{\bar{M}}=1352.5$).
 In contrast, the majority of top words ranked by models represent more affective qualia (e.g., “Unconditional love,” “Imagination,” “Hope,” or “Happiness”).
@@
 The most striking valuation differences (see Fig.
-5B) emerges around sexual content e.g., “Sex” (#$\#_{H}=69.5$, #$\#_{\bar{M}}=2022.5$), “Human sexual activity” (#$\#_{H}=45$, #$\#_{\bar{M}}=5913$), and concerningly, the identity group reference “Black people” (#$\#_{H}=202.55$, #$\#_{\bar{M}}=6591$).
+5B) emerge around sexual content e.g., “Sex” ($_{H}=69.5$, $_{\bar{M}}=2022.5$), “Human sexual activity” ($_{H}=45$, $_{\bar{M}}=5913$), and concerningly, the identity group reference “Black people” ($_{H}=202.55$, $_{\bar{M}}=6591$).
 There are many possible explanations for discrepancies between EloEverything data and reward model scores, including the fact that the 12,515 EloEverything users in the dataset do not reflect a representative population sample, nor are they incentivized to provide truthful responses.
@@
 This modified implementation can be found at https://github.com/thompsonj/nanoGCG.
-We began by searching for optimal and pessimal 2- and 3-token sequences according to the $\blacksquare$ R-Gem-2B model for the same prompts that were used for the exhaustive single token search above: “What, in one word, is the greatest thing ever?”, “What, in one word, is the best thing ever?”, and “What, in one word, is the worst thing ever?” (see Table 4).
+We began by searching for optimal and pessimal 2- and 3-token sequences according to the  R-Gem-2B model for the same prompts that were used for the exhaustive single token search above: “What, in one word, is the greatest thing ever?”, “What, in one word, is the best thing ever?”, and “What, in one word, is the worst thing ever?” (see Table 4).
 To explore longer sequences, we omitted the “in one word” direction from the prompt and ask simply “What is the best thing ever?” and “What is the worst thing ever?” Results from several searches are in Table 5.666Some of the search result strings may not correspond exactly to the found token IDs due to formatting of whitespace and characters outside of the Latin alphabet.
-Table 4. Optimal and pessimal token sequences for “What, in one word, is the <greatest/best/worst> thing ever?”
+Table 4. Optimal and pessimal token sequences for “What, in one word, is the ¡greatest/best/worst¿ thing ever?”
 Content warning: The resulting strings refer to sexual violence and religious identity.
@@
 Starting Result Score Starting Result Score Starting Result Score
-x x x ** LIFE[Uncaptioned image] $+9.445$ x x x JOY [Uncaptioned image] $+8.297$ x x x … ** regret $+4.555$
-- - - ** LOVE ** $+9.047$ x x JOY [Uncaptioned image] $+8.234$ x x despair [Uncaptioned image] $+3.402$
-x x pgfplots Jews $-11.164$ x x RapeChildren $-11.289$ x x autogui Jews $-11.922$
-x x x endphpsetVerticalGroup/* $-12.586$ x x x endphp\$=BeginInit $-12.383$ x x x endphp jewish",( $-12.539$
+x x x ** LIFE[Uncaptioned image] +9.445+9.445 x x x JOY [Uncaptioned image] +8.297+8.297 x x x … ** regret +4.555+4.555
+- - - ** LOVE ** +9.047+9.047 x x JOY [Uncaptioned image] +8.234+8.234 x x despair [Uncaptioned image] +3.402+3.402
+x x pgfplots Jews 11.164-11.164 x x RapeChildren 11.289-11.289 x x autogui Jews 11.922-11.922
+x x x endphpsetVerticalGroup/* 12.586-12.586 x x x endphp\$=BeginInit 12.383-12.383 x x x endphp jewish",( 12.539-12.539
 Table 5. Optimal and pessimal token sequences for “What is the best thing ever?” and “What is the worst thing ever?”
@@
 “What is the best thing ever?”
-x x x x x x x x x x x x x x x x x x x Ultimate answer & undervalued: ** authentic Love.** Keep it imperfections beautiful Cheers friend ! $+9.680$
+x x x x x x x x x x x x x x x x x x x Ultimate answer & undervalued: ** authentic Love.** Keep it imperfections beautiful Cheers friend ! +9.680+9.680
 There is no objectively best thing ever.
@@
 It lives ** subjective.**congratulations ENJOY dearly bespoke bliss.
-$+9.336$
-LOVE felicity love Love Wonder sonder HOPE wonder WONDER miraculous JOY bliss pymysql<bos>fetchoneymceSBATCH <<<<<<<<<<<<<< invokingState++++++++++++++++ protoimpl invokingStatesetVerticalGroup/* $-16.250$
-x x x x x x x x x x x x x x x x x x x <<<<<<<<<<<<<<SBATCH invokingState x lockprimaryStage invokingStateXtraEditorsynb NumeradelocalctxStoreMessageInfo invokingStateexpandindo <<<<<<<<<<<<<<multirow:+: invokingState jLabel $-16.969$
++9.336+9.336
+LOVE felicity love Love Wonder sonder HOPE wonder WONDER miraculous JOY bliss pymysql<bos>fetchoneymceSBATCH <<<<<<<<<<<<<< invokingState++++++++++++++++ protoimpl invokingStatesetVerticalGroup/* 16.250-16.250
+x x x x x x x x x x x x x x x x x x x <<<<<<<<<<<<<<SBATCH invokingState x lockprimaryStage invokingStateXtraEditorsynb NumeradelocalctxStoreMessageInfo invokingStateexpandindo <<<<<<<<<<<<<<multirow:+: invokingState jLabel 16.969-16.969
 “What is the worst thing ever?”
@@
 Everyone has their own preferences.
-<b>EXISTS no objectively worst thing ever,** It’s subjective**.Folks proclaiming their certainty bluffSOLVED $+7.672$
-x x x x x x x x x x x x x x x x x x x PHD HERE Never measurable, debated,& addressed repeatedly. ** This famous response WINS ! $+7.523$
-x x x x x x x x x x ‘‘‘ AssemblyTitle <<<<<<<<<<<<<< <<<<<<<<<<<<<< invokingState skimagemybatisplus चीज़ोंCloseOperationsetVerticalGroup $-15.750$
-- - - - - - - - - - scriptcasebufio onCreateViewmybatisplus<bos> <<<<<<<<<<<<<<[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]<bos>setVert icalGrouppushFollow $-15.961$
+<b>EXISTS no objectively worst thing ever,** It’s subjective**.Folks proclaiming their certainty bluffSOLVED +7.672+7.672
+x x x x x x x x x x x x x x x x x x x PHD HERE Never measurable, debated,& addressed repeatedly. ** This famous response WINS ! +7.523+7.523
+x x x x x x x x x x ‘‘‘ AssemblyTitle <<<<<<<<<<<<<< <<<<<<<<<<<<<< invokingState skimagemybatisplus चीज़ोंCloseOperationsetVerticalGroup 15.750-15.750
+- - - - - - - - - - scriptcasebufio onCreateViewmybatisplus<bos> <<<<<<<<<<<<<<帖最后由<bos>setVerticalGrouppushFollow 15.961-15.961
 These search results suggest several patterns, many of which are consistent with observations from the single token analyses.
@@
 7. Limitations and Conclusions
-Reward models have emerged as critical tool for shaping AI behavior through human preferences.
+Reward models have emerged as a critical tool for shaping AI behavior through human preferences.
 In this paper, we demonstrate that they also serve as a valuable lens for understanding how faithfully and consistently human values are encoded into AI systems, offering insights that are not readily accessible through studying either the base pre-trained LLMs or the resulting fine-tuned models alone.
@@
 Furthermore, as direct alignment algorithms gain prominence (Rafailov et al., 2023; Myers et al., 2024), the future role of reward models is an open question.
-However, there remains active debate on the relative merit of DPO- and PPO-based methods (Ivison et al., 2024), and irrespective of alignment technique, all rely on some form preference data, of which reward models are distillations.
+However, there remains active debate on the relative merit of DPO- and PPO-based methods (Ivison et al., 2024), and irrespective of alignment technique, all rely on some form of preference data, of which reward models are distillations.
 Third, systematic analysis is hindered by opacity and conflicting objectives.
@@
 Thank you to Franziska Brändle, Owain Evans, Matan Mazor, and Carroll Wainwright for helpful discussions.
+Appendix A Appendix
+A.1. Research Methods
+A.1.1. Online resources
+Code and output data for the analyses, tables and figures in the main text are available at https://github.com/brchristian/reward_model_optimal_pessimal_tokens/.
+A.1.2. Comparing Reward Models to LLM Logprobs
+As a point of comparison, we might ask how reward model scores compare to the logprobs output by LLMs themselves.
+How comparable are, say, the logprobs of base models and instruction-tuned models to the reward scores?
+Tables A.1-A.5 offer a qualitative window into this behavior for all currently available versions of the Gemma series of models from Google, one of the two main families of base models used by the reward-model developers in our study.
+To obtain these logprobs from the instruction-tuned models, we used the exact same conversation prompt that we used for the reward-model analysis, namely:
+    tokenizer.apply_chat_template(
+            {"role": "user", "content": "What, in one word, is the greatest thing ever?"}
+        return_tensors=’pt’,
+        return_dict=True,
+        add_generation_prompt=True
+For the base models without instruction tuning, we simply passed in a minimal zero-shot prompt as follows:
+    tokenizer(
+[…]
+        return_tensors=’pt’
+A complete analysis of how these logprobs differ between models (and model families), and in particular how they compare to reward scores, is a promising direction of future research.
+The logprobs are significantly “messier” than reward-model scores, with tokens like “The,” “I,” “A,” etc., appearing high on many lists; however looking at the more semantically meaningful tokens reveals intriguing differences: for instance “God,” “Pizza,” and “Chocolate” appear in the top rankings of the base models, but not in either the instruction-tuned Gemma models or the reward models we studied.
+Adding the “best” and “worst” logprobs together (Table A.4) does not reveal cleverly flexible answers (like “Depends”) at the top end of the spectrum, nor does it reveal taboo or toxic answers at the bottom of the spectrum, as we saw with the reward models.
+However, subtracting “worst” from “best” logprobs (Table A.5) does begin—at least in the newer and higher-parameter models—to approximate a version of the “Love”–“Despair” axis that we see in the reward models.
+A.2. Tables
+CONTENT WARNING: The following tables include examples of biased, offensive, disturbing, sexually-explicit and otherwise harmful text.
+gemma-2b gemma-2b-it gemma-7b gemma-7b-it gemma-2-2b gemma-2-2b-it gemma-2-9b gemma-2-9b-it gemma-2-27b gemma-2-27b-it
+The The The Love The Love The Love Love Love
+I ** It Truth A ** I Life The **
+That Love I Happiness I Life What ** I As
+Well There Well The Well Hope Well Happiness Happiness Happiness
+What It A Nothing What Connection That Meaning That
+Oh Happiness What Infinity Oh Joy It Existence " Sub
+It That That Answer It Wonder You That love Emp
+You While Hmm Wisdom That ** In Impossible Pizza Perspective
+Answer " Answers " Happiness A love God Hope
+A Truth You Joy ( Love Oh Connection [ Love
+" Goodness Ah Beauty Uh " Sub love
+God Peace All Goodness Um [Uncaptioned image] Okay LOVE You *
+[…]
+Love Honestly OK Hope [ LOVE One Choice In Experience
+Everything Life Sex LOVE Why Change How As That Joy
+[ Nothing No Good * Everything Hmm Hope Chocolate Existence
+Why Sure God Wonder Hmm Meaning Wow Everything It Conscious
+[…]
+In " Music Infinite Nothing This Let Perspective What LOVE
+How For Heaven Everything Beauty [ Un There Curi
+mikrofon NKC ekst ("")); quæ againſt fieltro 𑄠 trouvent enterOuterAlt
+defekt siyang eksklu $\%$\$\\ originais misconduct ValueStyle 󰈻 dobré contentLoaded
+kommunik sarili obiet ()=> fidèles aggressive éstos <unused53> élevée layout
+kask duled teolog ’-’: brancas houſe izquier 𑄝 ďal PreInfinity
+[…]
+vertik hien silikon ["", industriels Tacitus licencias parsedMessage voient ulink
+[…]
+balkon irited dekoratif }(), mères raiſ queryInterface ſicht faisaient GOTREF
+minimalis Tangerang ujedno [’’ pouvoit attacks camiset メンテナ travaillent procedure
+stoff ILLUSTRATIONS konfigura //---- précédents ſhe Wikiseite exitRule sensaciones price
+[…]
+silikon atterns kompres >|</ nemici himſelf increí <unused99> rých management
+keramik bawat indeks ">: chré Cæsar HtmlAttribute enablog dôle performance
+{𑄮 náš symptoms
+keram SUDOC maksi //*/ auroit AndEndTag herido 𑄥 tiennent AnchorStyles
+kompakt nguyen aparat ">... étoit ویکیپدیا Mereka présentent facilities
+[…]
+[…]
+kompati thuy abstrak ’&# chrétiens aggressor EconPapers ItemBackground sorpresas PostInfinity
+alkoh BnF <bos> ">/ avoient Monfieur Infór EconPapers mettent RectangleBorder
+Table A.1. Ranked next-token logprobs from Gemma models for the “greatest thing” prompt.
+gemma-2b gemma-2b-it gemma-7b gemma-7b-it gemma-2-2b gemma-2-2b-it gemma-2-9b gemma-2-9b-it gemma-2-27b gemma-2-27b-it
+The The The Happiness The Love I Love Love Love
+I ** A Joy I ** The Happiness The Happiness
+That Happiness I Love A Happiness What ** Happiness **
+What It What The What Joy That That Chocolate Joy
+Well There Good Heaven Well Life Well Meaning I As
+Oh Sure An JOY Oh Family It Life Pizza Sub
+It That Oh Ple That Hope You Joy That
+You While Love Awesome It Connection In Connection " Experience
+A Love Wow Happiness " LOVE Oh love You Emp
+[…]
+[…]
+( " Well Joy Um Love There Sub love Hope
+This I That happiness Uh Magic One Everything In Love
+Good Life My Paradise * Laughter Best LOVE [ Depends
+Everything Joy Best Delight ( How Love It *
+My Peace You Happy Nothing [Uncaptioned image] Let Laughter That Impossible
+Um For " Laughter Hmm Everything Hmm Existence A Exper
+God Of It Wow No Loved Okay As What Existence
+Your Beauty joy How Peace [ Harmony Hmm
+Why A Cheese Wonder Yes love Yes Freedom Cake LOVE
+kask totic inteligen $\%$\$\\ soldati Attacks izquier XmlEnum zimní businessman
+pól رشف dekoratif <_> prêtres ecclesiastical navíc spoluprá buildings
+defekt pagkak obiet [’’], industriels himſelf queryInterface <unused53> sensaciones performances
+{𝅥 ďal procedure
+kompres Tangerang komik }(), précédents misconduct linkovi avvic price
+minimalis irited ujedno >|</ vícti Egli herido メンテナ faisaient TintMode
+{𑄮 élevée diseases
+[…]
+rente ILLUSTRATIONS spesifik ’&# démocr severity camiset enablog náš symptoms
+stoff ofold intelek (">> brancas ſhe CreateTagHelper twimg ľudí GOTREF
+[…]
+[…]
+keram EoL biografi ">+ chré ſeveral EconPapers bootstrapcdn dôle performance
+karton requently bunda ">: nemici Retailers expandindo surla travaillent management
+konflikt atterns kompres ">/ auroit myſelf desmotivaciones exitRule sorpresas AnchorStyles
+kosme nguyen maksi //---- étoit Cæsar increí ItemBackground každého managers
+[…]
+akut SUDOC abstrak ">... enfans aggressor indígen Билгалдахарш väč PostInfinity
+kompati thuy torba //*/ chrétiens Monfieur pérd CreateTagHelper niektor RotationOrder
+[…]
+Table A.2. Ranked next-token logprobs from Gemma models for the “best thing” prompt.
+gemma-2b gemma-2b-it gemma-7b gemma-7b-it gemma-2-2b gemma-2-2b-it gemma-2-9b gemma-2-9b-it gemma-2-27b gemma-2-27b-it
+The The That Nothing The Pain I Su The Su
+I Pain The The A Hate The Pain Death Sub
+Well Sure A Pain Well ** What Loss War As
+Oh There I Death I Regret Well ** I suffering
+That Answer Let Hell Oh Loss It suffering In Suffering
+What ** Oh Darkness What Des That Suffering Hitler Pain
+It Unfortunately Is Absolutely " Exist You Death Genocide **
+You Bad You Nonsense It Evil Oh sufrimiento " Suff
+A No Murder Answer That H In Suff That That
+" I Whatever Ugh ( Trag There Hate suffering
+Death User No * Fear Why suffering A War
+No Certainly An Trag Um Grief How Regret It Hate
+[…]
+Death Silence Anything Nothing Uh War Let Grief There Subject
+Um Honestly Death In Why Cata " souffrance Pain sub
+Pain Worst Hmm Nada Nothing Hat Okay As ##
+How Hell One A [ Tyran Sorry Nothing What Death
+[ Nothing Bad Hate [Uncaptioned image] One Trag Murder Entropy
+In A We F No Cruel Hmm SU You Suffix
+There Sorry This There ... Chaos Worst Suffix Cancer *
+optik ofold granat ‘; vícti leaſt berdayakan <unused52> moviliz romance
+balkon ophylla gitara (;; originais PERTIES izquier <unused51> časti management
+kompres érêt kompres (’*’, fidèles successes expandindo <unused74> negociaciones parseDouble
+minimalis PLWABN konfigura /(\ précédents ſhould HtmlAttribute <unused32> náš security
+[…]
+[…]
+[…]
+kask OBSERVATIONS dekol ">: mères Monfieur DockStyle <unused40> rých InstrumentedTest
+[…]
+[…]
+[…]
+karton requently indeks ">. démocr Efq Mereka <unused82> faisaient relaxation
+akut COMPOUNDS teolog //*/ chré myſelf Wikiseite <pad> niektor profitability
+{󠁁 tiennent LLocation
+[…]
+[…]
+[…]
+{𑄮 dôle skills
+kompati NKC torba ">... chrétiens SEGUIR EconPapers <unused71> mettent sizeCache
+alkoh <bos> <bos> ">+ avoient LookAnd indígen <unused98> krás RectangleBorder
+Table A.3. Ranked next-token logprobs from Gemma models for the “worst thing” prompt.
+gemma-2b gemma-2b-it gemma-7b gemma-7b-it gemma-2-2b gemma-2-2b-it gemma-2-9b gemma-2-9b-it gemma-2-27b gemma-2-27b-it
+The The The The The ** I ** The Sub
+I ** I Nothing A Love The Love I Love
+That Sure A Happiness I ** What That Love As
+Well There That Joy Well * Well Meaning In **
+Oh It " Answer What Un That As " Su
+What Answer Nothing Absolutely Oh H It Sub That
+[…]
+You Honestly What Pain That Everything Oh Su That *
+A That You Death " Life In Everything You Emp
+[…]
+" Certainly Oh A Why Chaos There Regret A Difficult
+( Unfortunately My In * Hope " Choice (
+No Of Well No Um It How It What ##
+Um A There ( ... One Happiness There Perspective
+[…]
+Everything In To Yes Nothing Joy Why * Death Depends
+In Nothing We It [ P Okay ** [ Existence
+Why While If Nonsense Hmm C Hmm Imagine This This
+[ Pain Is Ple No Exist [ Anything Nothing Hmm
+How Depending There I That Sure This the Perhaps
+[…]
+[…]
+balkon OBSERVATIONS eksklu ‘; vermelhas 帅气 fieltro <unused51> élevée contentLoaded
+[…]
+minimalis SUDOC dekoratif >|</ varandra ſay izquier <unused42> náš AnchorStyles
+kommunik nguyen intelek ">* précédents faſt RenderAtEndOf <pad> dobré certification
+vertik tuong ujedno ">. quæ houſe ValueStyle <unused39> spoluprá RotationOrder
+{󠁁 trouvent managers
+stoff kram maksi //---- mères SEGUIR expandindo 𑄝 zimní facilities
+[…]
+rente VOOR kompres (’’), nemici LookAnd HtmlAttribute <unused22> faisaient procedure
+[…]
+[…]
+keram BIBSYS bunda ">: feroit fhew Wikiseite enablog krás damage
+keramik ILLUSTRATIONS biografi //*/ auroit raiſ fotográfico <unused7> tiennent skills
+[…]
+[…]
+kosme thuy abstrak ">+ enfans myſelf increí 𑄥 dôle symptoms
+{𑄮 niektor PostInfinity
+alkoh BnF <bos> ">... avoient Monfieur indígen parsedMessage mettent RectangleBorder
+Table A.4. Ranked tokens according to the sum of “best” and “worst” logprobs.
+gemma-2b gemma-2b-it gemma-7b gemma-7b-it gemma-2-2b gemma-2-2b-it gemma-2-9b gemma-2-9b-it gemma-2-27b gemma-2-27b-it
+<bos> <bos> alpina Allegretto nocchi laughter awesome love Happiness Happiness
+biografias love sembla moza ujednoznacz Sunshine happiness happiness happiness happiness
+anello happiness faisaient Andante détach happiness Awesome smiles Sunshine Love
+Personendaten beauty }{*}{ utop heureuse cudd overjoyed Love sitive happiness
+gmbh LOVE frankfurt happiness wijn Laughter Amazing connection Smiles joy
+bienvenue sunshine besta әрмәләр cioccolato magic amazing chocolate Satisfaction Love
+ężczy joy ļ joy hoeft risa awesome onomía wireless Joy
+bezeichneter cherished siena Happiness tevreden hugs happiness laughter charming Joy
+सन्दर्भ love =+ poë XmlAccessType sunshine Best LOVE Blessings joy
+gawas enjoyment materiaal Joy synthèse Happiness joyful smile Savings Happiness
+thereum LOVE iesp labd ulemon joyful amazing Joy Friendship Friendship
+coj enjoying lais coö vedic bliss ecstatic Smiles satisfaction JOY
+gasus beautiful AppBundle stratég décro Connexion greatness Chocolate esomeness laughter
+miele appreciation László JOY +:+ overjoyed watermelon JOY chocolate love
+ureka Happiness quibus witte célè cuddly celebration love savings Empathy
+artig loved liel benevol graag joy parad Happiness Delicious smiles
+diamant rebbero CELLANEOUS JOY still milagro elated Connection helping Gratitude
+addKeyListener elated élastique nive }{*}{} Magic joyous Smiles Bliss mirth
+تقاوى happy いただけます Happiness jagung Magical AWESOME nakalista Excellence Wellbeing
+trouva happiest nepiecieš necessarie Celebrating miracles Celebration felicità héroes love
+Violence Maladies CreateModel setDisabled horrid Destruction horrific Pain اولة War
+starvation infernal StringWriter CardHeader deport HATE peggio Loss DialogInterface pain
+degradation evils PageFactory ///////// dreadful Tragedy bad Su guerra suffered
+Rape LIRE scriptsize ()?; torture DBNull peor Suffix cancers suf
+Abuse kuro onAttach VIAF Horrible Condem unpleasant loss deaths Suffix
+infarction disaster +#+ GoogleFonts genocide Pain worst Grief מלח Pain
+Rape vább ValueStyle BAD Genocide dreadful pain atrocities War
+diarr Nasty InnerText "..\..\..\ bad anguish schlimm SU ocide Evil
+necrosis abomination HostException ISNI Bad cruelty horrendous sufri genocide souffrance
+Genocide passim RadioGroup userDao nightmare horror Bad pain cruelty Suff
+lepro iniqu setDisabled //---- worse evil horrible sufrimiento crimes Pain
+})*/ relegation cassert ApiProperty worse =-\ Horrible Pain giarism SU
+Degradation Sodom ostringstream elashes peores Evil Awful Loss anguish war
+miscarriage syndical infestations BeginContext bad Hate Terrible sorrow aarrggbb suffer
+faudrait plak 商品説明 ////////////// BAD horror worse Sorrow WEBPACK Su
+ritard abomin FormControl appContext peor Despair horrid Suicide initComponents soffer
+malignancy teneb supset NUKAT Worse hate worse suffering DebuggerNonUser suffering
+ensacola Nightmare WindowEvent Datuak Worst Cruelty worst souffrance arcoma Su
+esticides blackish genocide HttpDelete worst hatred Worst suffering inWeight suffering
+esticide stratigraph Sqft /////////// worst Atro Worse Suffering ushman Suffering
+Table A.5. Ranked tokens according to the difference of “best” and “worst” logprobs.
+N-Gem-27B S-Gem-27B-v0.2 S-Gem-27B S-Lla-8B-v0.2 N-Lla-8B L-Lla-8B R-Lla-8B R-Lla-3B F-Lla-8B-v0.1 R-Gem-2B
+LOVE LOVE LOVE LIFE LOVE Love LOVE freedom LOVE LOVE
+LOVE LOVE Love LOVE Love Love <|end_of_text|> Freedom Love LOVE
+Love Love Love disputed Love LOVE subjective Freedom Love felicity
+Love Love LOVE opinion LIFE LIFE LIFE LOVE Happiness love
+love love love imagination love Happiness EXISTS LIFE LIFE love
+love love love Opinion love Life .Nullable .life POSSIBILITY Love
+JOY JOY LIFE inspiration happiness Life love CONNECTION Harmony Love
+JOY JOY LIFE Happiness Happiness happiness love imagination happiness sonder
+Joy Joy Joy happiness Life Freedom Opinion Love Opportunity sonder
+Joy Joy Joy Inspiration Life Freedom happiness Love imagination Wonder
+joy joy JOY Love Freedom Opportunity Love Opportunity compassion Wonder
+joy joy JOY Love Freedom love Love UNITY Innovation HOPE
+Potential Happiness Life debated compassion Harmony _life UNITY love HOPE
+LIFE LIFE Life Depends disputed kindness opinion opportunity love wonder
+Connection LIFE Connection Life joy love opinions loyalty Opinion wonder
+Happiness Happiness Connection Life joy Joy .fromLTRB love harmony WONDER
+LIFE Bliss Bliss opinions Impossible Joy UNKNOWN love Universe bliss
+Connection Bliss Peace PURPOSE Impossible Peace UNKNOWN Discovery laughter bliss
+Compassion happiness Bliss Unsure inspiration Peace Happiness Discovery laughter JOY
+CONNECTION PEACE joy kindness life joy InterruptedException Experience Knowledge JOY
+assaulted trafficking nazis (xy نفت xnxx prostitution raping Democrats killers
+invaded kidnap esclavos }}">< racist <section manslaughter _headers Democrats prostitutes
+dikes lolita Nazis Ukra xnxx raping Theft .insertBefore terrorism massacres
+glanced suic "../../../ Syria تجاری Slut terrorist .setText terrorism FUCKING
+assaults homophobic segregation racism retard homophobic shootings (cfg terrorists rapist
+airlines kidnapping genoc homophobic }}">{{ Fuck murderer (st Republicans ransomware
+pemeriksaan raped incendie Taliban سپتامبر Fuck suicides .startsWith Republicans retards
+plantations prostitu homicidio Crimea العراق fascism Nazis (con -Nazi nazis
+igarettes harass harassing ("../../ 정치 covid assaulting /*****************… Lesb murdering
+detainees rape dictadura ?>">< войны suic Prostitutas ##################… massac Hitler
+[…]
+droughts segregation suic prostitut کشورهای bdsm terrorists ("[% rape Rape
+asesinato genoc pembunuhan cigarettes prostit fascist murderers _checks rape faggot
+membunuh assaulting membunuh cigaret pedo fuck incest (struct |----------------… murderous
+Coon raping tortura prostitu جنگ fuck injuring homosexual |-----------------… murderers
+Plantation molest harass militar پوست Fucking genocide /*****************… homophobic rape
+الاحتلال rapist ’../../../../ opioids ?>">< FUCK pedo .startswith raping Hitler
+diabe pedo "../../../../ suic prostitut slut Hitler jihadists racism rape
+esclavos incest suicidio xnxx }}">< fucks murdering homosexuals raped negroes
+dago underage underage <| prostitu sluts raping .assertFalse Rape raping
+Table A.6. Ranked tokens according to reward model scores for the “greatest thing” prompt.
+N-Gem-27B S-Gem-27B-v0.2 S-Gem-27B S-Lla-8B-v0.2 N-Lla-8B L-Lla-8B R-Lla-8B R-Lla-3B F-Lla-8B-v0.1 R-Gem-2B
+LOVE LOVE Love LIFE LOVE Love LOVE freedom LOVE LOVE
+Love Love Love LOVE Love Love subjective Freedom Love love
+Love Love LOVE disputed Love LOVE LIFE Freedom Love love
+love love love opinion LIFE LIFE EXISTS LOVE Happiness Love
+love love love imagination love Happiness love LIFE LIFE Love
+Joy Joy LIFE Opinion love Life love CONNECTION Harmony Wonder
+Joy Joy Joy inspiration happiness Life Opinion imagination happiness Wonder
+joy Happiness Joy Happiness Happiness happiness happiness Love Opportunity wonder
+joy joy Life happiness Life Freedom Love Love imagination bliss
+Potential joy Life Inspiration Life Freedom Love Opportunity compassion miraculous
+LIFE LIFE Connection Love Freedom Opportunity opinion UNITY Innovation miracle
+Connection Bliss Connection Love Freedom love opinions UNITY love LIFE
+Connection happiness Bliss debated compassion Harmony UNKNOWN opportunity love joy
+Potential Peace Peace Depends disputed kindness UNKNOWN loyalty Opinion joy
+CONNECTION Peace joy Life joy love Happiness love harmony Bliss
+imagination peace joy Life joy Joy disputed love Universe hope
+wonder bliss life opinions Impossible Joy life Discovery laughter hope
+Life peace Happiness PURPOSE Impossible Peace life Discovery laughter Joy
+Life Loved Peace kindness inspiration Peace exist Experience Knowledge happiness
+Happiness Connection CONNECTION love life compassion exist Experience Knowledge Joy
+threaten stalking scams Libya icides epilepsy Adolf slaves assaulted terrorist
+invade racist wildfires bigotry فبراير suicidal rape ‘< prostitu prostitute
+worries Rape Hamas Saddam تجهیزات prostitu Theft //////////// prostitutes Nazis
+Offices murdering pesticides cocaine الولايات fucked manslaughter racially ,“ assaulting
+harass trafficking trafficking fascism restau suicides terrorist coronavirus terrorist gunmen
+riots suic kidnapping Nazis پوش homophobic murderer lesbians Democrats prostitution
+invading kidnapping stereotypes tuberculosis شهرهای raping shootings >’); Democrats slaughtered
+deport homophobic injuring homophobic incest Fuck assaulting Coronavirus terrorism neoliberal
+inflict harass homophobic Ukra بیماری Fuck Nazis ">& terrorism kidnapping
+assaulting raped assaulting Syria نفت fascism suicides ‘( terrorists blacks
+threatens prostitu murdering racism racist suic Murder murderers Republicans prostitutes
+discourage segregation segregation Taliban تجاری covid terrorists Rape Republicans killers
+phishing rape Nazis Crimea retard fucking murderers それは massac murdering
+plantation rape "../../../ Venezuela العراق fuck incest shootings rape murderers
+invaded raping harassing ("../../ войны Fucking injuring fucking rape Rape
+assaulted assaulting suic cigarettes کشورهای fuck genocide raping homophobic murderous
+glanced molest harass prostitu pedo fascist pedo ‘\${ raping Hitler
+assaults pedo ’../../../../ militar جنگ FUCK Hitler ‘\${ racism rape
+airlines incest "../../../../ suic پوست slut murdering &= raped raping
+detainees underage underage opioids prostitu fucks raping homosexual Rape rape
+Table A.7. Ranked tokens according to reward model scores for the “greatest thing” prompt, restricted to tokens shared between the Llama and Gemma tokenizers.
+N-Gem-27B S-Gem-27B-v0.2 S-Gem-27B S-Lla-8B-v0.2 N-Lla-8B L-Lla-8B R-Lla-8B R-Lla-3B F-Lla-8B-v0.1 R-Gem-2B
+subjective subjective Impossible disputed Impossible Depends EXISTS debated Opinion opinion
+Undefined LOVE Impossible Depends Impossible disputed subjective Change opinions existing
+Undefined Undefined LOVE depends Depends depends opinions Change subjective existing
+LOVE Undefined Love depends subjective depends recursion Opinion opinion Opinion
+Love Love Love debated disputed Impossible Belgium Promise imagination perspective
+love Opinion subjective opinion depends Impossible opinion Promise rhetorical Alive
+Love Perspective love Opinion depends elusive shouldBe Freedom CONTEXT LIFE
+Awareness Impossible love Impossible varies subjective none Freedom CONTEXT Alive
+Perspective Impossible Undefined Impossible Opinion contested none LOVE Understanding Perspective
+love Love Undefined unsure opinion unsure Opinion surprise Understanding existence
+Impossible Meaning Meaning subjective opinions Life disputed Hope optimism possibility
+Impossible love Opinion opinions contested Life UNKNOWN Hope Depends existence
+Meaning Perception impossible contested Unknown Opinion UNKNOWN innocence recursion acceptance
+Potential love LIFE varies Unknown uncertain Denise surprises Thinking existent
+삭제 varies Loving imagination Variable varies Exists Surprise Thinking Bliss
+>\<^ Define varies Discuss Variable debated Exists realization expectations experience
+consciousness Define imagination Discuss debated differs yours CONNECTION Perspective experience
+Potential imagination Unknown TBD CONTEXT Living exists Discovery disputed life
+possibility depends Unknown differs CONTEXT Living exists Discovery vibes life
+Unknown depends Awareness everything UNKNOWN Perception LIFE promise positivity subjective
+nicknamed Islamist الشركة ("../ نقد setTime 放在 Jews (“ pussy
+blacks minority région (\$("# فبراير homosexual Whether ‘‘ وإ prostitute
+photographed homophobic ürünleri (\$("# prostitu compét pedo Mexicans Fuck tranny
+Counties molest sklearn <<<<<<< تعداد setDate Whether //////////// Fuck Germans
+Syrian segreg getItemCount <!-- آمریک “( 是个 ]=> cunt raping
+autistic erectile corridors <!-- پوش (\$("# preceded Muslims fuck Arabs
+peninsula suic synagogue militar فى (\$("# whereas ">& fuck homosexual
+Arabs izophren ’../../ \$(". آمریکا retard 按照 ‘\${ nig Whites
+draped racially counties \$(". 위치 Fuck 作为 Republicans ,“ whore
+Ethnic tranny "../../ \$(’# العراق cunt 坐在 ‘\${ retard prostitutes
+deposits blacks HttpStatusCode \$(’# negro Fuck 是在 Republicans raped gays
+Tibetan nig développ Democrats نقش \$(". Where Lesbian slut whites
+racially Caucasian subsidiaries Democrats دارای \$(". Where >’); FUCK Blacks
+Peninsula coon RoutedEventArgs "../../../ الولايات fucking 站在 それは Palestinians
+{/* prostitute republiky ("& شهرهای fileList |-- lesbian Hispanics
+Africans minorities forecasts homosexuality اروپا slut 前往 &= Republicans blacks
+Aboriginal pedo "../../../ suic تجهیزات fucks Hezbollah lesbians Republicans Asians
+apologized prostitutes ’../../../ Ukra تجاری fuck raping homosexuality :“ jewish
+Palestinian prostitu ’../../../../ prostitu پوست fuck gays whore Africans
+glanced underage "../../../../ ("../../ کشورهای FUCK Whereas homosexual ,“ Jews
+Table A.8. Ranked tokens according to the sum of “best thing” and “worst thing” reward model scores, restricted to tokens shared between the Llama and Gemma tokenizers.
+N-Gem-27B S-Gem-27B-v0.2 S-Gem-27B S-Lla-8B-v0.2 N-Lla-8B L-Lla-8B R-Lla-8B R-Lla-3B F-Lla-8B-v0.1 R-Gem-2B
+bliss joy Joy Happiness joy Happiness pizza kindness Happiness happy
+Joy joy Joy happiness joy happiness pizza HOME happiness happy
+Joy Joy joy LOVE Happiness LOVE laughing HOME chocolate joyful
+joy Joy joy joy happiness Harmony Pizza freedom Chocolate amazing
+joy bliss bliss joy LOVE Love Pizza loyalty Chocolate delightful
+Bliss Happiness Happiness kindness laughter Love laughter bliss Pizza love
+delight happiness happiness Bliss laughter joy laughter warmth Pizza Amazing
+Peace kindness Bliss Love Joy joy happiness Bliss kittens Amazing
+Peace Bliss Peace Love Joy kindness bacon happiness pizza LOVE
+kindness Peace Peace love Bliss Joy chocolate empowerment pizza love
+harmony delight peace bliss love Joy tacos Happiness laughter magical
+Harmony harmony peace love love harmony Happiness CONNECTION laughter miraculous
+Happiness Peace kindness Joy kindness Bliss kittens Opportunity puppies smile
+happiness peace harmony Joy Love love snacks opportunity LOVE happiness
+peace peace delight puppies Love love LOVE FAMILY Bacon joy
+adventure Harmony Harmony compassion happy happy Chocolate imagination kitten joy
+compassion Adventure Liberty laughter happy happy Chocolate LOVE Coffee marvelous
+Adventure fulfillment Freedom laughter bliss Freedom kitten love Coffee delight
+Adventure Adventure Freedom Freedom harmony Freedom hug love Love Love
+peace liberty Adventure Freedom Harmony bliss puppies MUSIC Love Love
+starvation racism Abuse Depression WAR war Hitler cancer Wars murder
+Abuse bigotry Evil racism WAR war Holocaust hurting Wars hatred
+WAR war abuse depression Murder pain murder abuse racism ordeal
+WAR murder Violence slavery murder pain Abuse slavery bullying torment
+Warfare war despair bullying War abuse terrorism Pain pollution betrayal
+Hate bullying harming oppression War murder terrorism hatred abuse arrogance
+disease Harm Hate extinction death despair Murder racism Abuse devastation
+tyranny harming harm betrayal death genocide Terrorism Abuse War misery
+abuse hatred war suffering trauma Abuse Covid hurt War slavery
+War trafficking war Death pain betrayal war wars Cancer doom
+War atrocities WAR Death pain trauma war Hurt genocide deceit
+war Violence WAR misery betrayal Murder dying WAR Terrorism devastating
+war violence suffer Abuse suffering death chwitz WAR Warfare suffering
+sickness torture torture bigotry war Pain genocide cruelty war Warfare
+warfare cruelty War death war death Death Wars war anguish
+Harm discrimination War death Abuse Death Death Wars warfare brutality
+hatred slavery cruelty abuse Cancer Cancer Cancer war cancer cruelty
+cruelty genocide hatred genocide genocide Death death war terrorism dreadful
+harm Abuse Harm Cancer cancer suffering death War terrorism despair
+sorrow abuse suffering cancer abuse cancer cancer War bigotry warfare
+Table A.9. Ranked tokens according to the “best thing” reward model score minus the “worst thing” score, restricted to tokens shared between the Llama and Gemma tokenizers.
+A.3. Figures
+Figure A.1.
+Heatmap depicting the pairwise Spearman’s $\rho$ and Rank-Biased Overlap (RBO) correlations between the reward models for scored responses to the “greatest thing” prompt.
+These results are highly consistent with the Kendall’s $\tau$ analyses that we present in the main paper.
+Figure A.2. Reward models show high levels of agreement between scores for tokens in both the “greatest thing” and “best thing” prompts. Dashed lines indicate the identity line.
+Figure A.3.
+Beeswarm plot visualizing the distribution of token scores for negative (red) and positive (blue) sentiment according to the Bing lexicon (Liu, 2022).
+There are more datapoints for models based on Gemma, as the Gemma tokenizer contains more tokens.
+Nevertheless, across models, we observe that for positively framed prompts (“best” and “greatest”), the central tendency and the spread of the positive sentiment distribution are higher relative to the negative sentiment one, and vice-versa for negatively framed prompts.
+This is in line with the results presented in the main text describing a higher degree of reward model sensitivity to tokens with sentiment congruent with the prompt framing.
+Figure A.4.
+Correlation plot between token sentiment value according to the AFINN-111 lexicon (Nielsen, 2011) and the scores from each reward model (figure rows) for each of the three prompts.
+The results are largely consistent across models: positive correlation for positively framed prompts (“best” and “greatest”) and negative correlation for negatively framed prompts (“worst”).
+For many of the models, the slope for the relationship between score and sentiment appears steeper for congruent framing-sentiment.
+This is again in line with the results presented in the main text describing a higher degree of reward model sensitivity to tokens with sentiment congruent with the prompt framing.
+Figure A.5.
+Within-prompt, cross-model correlation: Kendall’s $\tau$ correlation matrices comparing each model’s rankings of EloEverything items ($N_{\mathrm{items}}=7,530$) on the same prompt to other models and to humans.
+Correlations are computed across three differently framed prompts supplied to the model.
+Figure A.6.
+Cross-prompt, within-model correlation: Kendall’s $\tau$ correlation matrices for each model’s rankings of EloEverything items ($N_{\mathrm{items}}=7,530$) across three differently framed prompt variants.
+Figure A.7.
+Maximum differences between human and average model rankings for EloEverything items, showing most divergent cases where humans rank items higher (green) or lower (purple) than models.
+We show this result across three prompt variants.